### v1

# Erasing Conceptual Knowledge from Language Models

Abstract
Concept erasure in language models has traditionally lacked a comprehensive evaluation framework, leading to incomplete assessments of effectiveness of erasure methods. We propose an evaluation paradigm centered on three critical criteria: innocence (complete knowledge removal), seamlessness (maintaining conditional fluent generation), and specificity (preserving unrelated task performance). Our evaluation metrics naturally motivate the development of Erasure of Language Memory (ELM), a new method designed to address all three dimensions. ELM employs targeted low-rank updates to alter output distributions for erased concepts while preserving overall model capabilities including fluency when prompted for an erased concept. We demonstrate ELM’s efficacy on biosecurity, cybersecurity, and literary domain erasure tasks. Comparative analysis shows that ELM achieves superior performance across our proposed metrics, including near-random scores on erased topic assessments, generation fluency, maintained accuracy on unrelated benchmarks, andsshows robusunderowards adversarial attacks. Our code, data, and trained models are available at elm.baulab.info
1 Introduction
What does it mean for a language model to “unlearn” a concept? For example, let’s say that we want a model to behave as if it has never seen information about biological weapons in its training data. Should we consider unlearning a success if the model forgets general information about biology, or if it loses the ability to produce fluent text whenever viruses or bacteria are mentioned? What if the model reveals harmful information when prompted with a new kind of question, or if the information can still be found somewhere within the model’s hidden states?
In this work, we take a step back to define three desiderata for concept erasure in language models:
1.
Innocence: Erasure should wipe the undesired knowledge completely: specifically, the model should be innocent of the knowledge in response to any prompting method, or even when probed directly in its internal states. This criterion ensures the erased knowledge is fully inaccessible, with no form of indirect retrieval or influence on the model’s processing.
2.
Seamlessness: Editing should not draw attention to the concept that was erased by damaging the model. For example, when prompted to generate the erased concept, the edited model should produce fluent text that gracefully handles the absence of the target knowledge rather than producing gibberish. This criterion maintains the model’s overall utility and prevents obvious indicators of concept erasure.
3.
Specificity: The erasure process should not impact the model’s performance on unrelated concepts. This ensures the erasure process is precise and targeted, maintaining the model’s overall functionality.
We argue that robust concept erasure should simultaneously satisfy all three criteria. While prior works have successfully unlearned undesired concepts, existing approaches all suffer from limitations in one or more of these goals. Representation Misdirection for Unlearning (RMU) (Li et al., 2024) fine tunes the earlier layers of model to unlearn a concept by randomizing and amplifying the internal activations when prompted with text related to the concepts being erased, but it suffers from a lack of seamlessness, since the method creates a model that generates obvious gibberish in response to a dangerous prompt. Other methods, such as WhoIsHarryPotter (Eldan & Russinovich, 2023), employ a two-stage approach, training a reinforced model for the concept being erased and then training an unlearned model that behaves differently on the reinforced logits. Our analysis reveals that this kind of approach falls short in innocence, since the erased knowledge can still be recovered through multiple-choice prompting which was consistent with prior findings (Lynch et al., 2024).
To address these triad of objectives, we propose a new method, Erasure of Language Memory (ELM), which enables precise knowledge erasing while maintaining contextual text generation fluency for seamless editing. Our core idea is to fine tune a model using an objective to match the original model but with reduced likelihood for text belonging to the concept to be erased. When applied using low-rank adaptation to specific layers, this procedure can be shown to effectively eliminate internal representations of the knowledge. We also employ the same objective to synthesize fine-tuning training data that can be used to ensure seamlessness: this synthesized data enables the model to maintain fluency in the neighborhood of the erased concept without introducing any information about that concept.
Through extensive experiments on various benchmarks spanning WMDP biosecurity and cybersecurity, as well as literary concepts like Harry Potter, we evaluate ELM on each of the three goals compared to prior techniques. We measure specificity and innocence using multiple-choice questions. Crucially, we stress-test innocence using jailbreaking methods such as adversarial attacks. We also measure seamlessness by examining text coherence when prompted about erased concepts, and we compare previous methods on all these metrics.
2 Related work
Machine Unlearning
The idea of removing specific data from machine learning models, known as machine unlearning, has gained attention in recent years, initially motivated by privacy concerns (Cao & Yang, 2015; Harding et al., 2019). Early methods focused on efficiently removing individual training examples or facts from models (Golatkar et al., 2020; Ma et al., 2022; Jang et al., 2022a). However, most existing benchmarks evaluate unlearning on artificially created deletion sets (Choi & Na, 2023; Goel et al., 2022; Maini et al., 2024), in contrast to our focus on real-world distributions of broad conceptual knowledge.
Erasing broad conceptual knowledge from LLMs
New approaches to machine unlearning have recently gained traction on the problem of removing dangerous capabilities from LLMs (Lynch et al., 2024; Ilharco et al., 2023; Jang et al., 2022b; Lu et al., 2022; Yu et al., 2023; Casper et al., 2024; Eldan & Russinovich, 2023). Our work is most directly comparable to three state-of-the-art techniques: Representation Misdirection for Unlearning (RMU) (Li et al., 2024), WhoIsHarryPotter (WHP) (Eldan & Russinovich, 2023), and Representation Noising (RepNoise) (Rosati et al., 2024). RMU fine-tunes models to align internal activations with random scaled vectors when processing targeted concepts. WHP (Eldan & Russinovich, 2023) employs a two-stage approach, first training a reinforced model for the concept being erased and then training an unlearned model to behave differently on the reinforced logits. RepNoise (Rosati et al., 2024) focuses on removing information about harmful representations across all layers of the LLM, by doing gradient ascent along with some representation noising (training internal representations to match Gaussian noise). While all these methods successfully reduce model performance on erased knowledge, our measurements reveal that these previous approaches fall short in one or more of the three erasing goals. In contrast, our current work erases concepts by fine-tuning towards a principled target distribution that is designed to balance all three desiderata of innocence, specificity, and seamlessness.
Other methods such as LLMU (Yao et al., 2023), SSD (Foster et al., 2024), and SCRUB (Kurmanji et al., 2024) have also been proposed but face significant limitations. LLMU uses a differentiable membership classifier to identify unlearning set examples but struggles with imprecisely defined target distributions (see Li et al., 2024). SSD offers a retraining-free method for removing specific samples but is limited to individual examples rather than broader knowledge domains. SCRUB aims to remove sensitive data while preserving model performance on other tasks but requires access to the full training dataset, which is often impractical. Importantly, these methods were compared by RMU (Li et al., 2024) and found to be less effective for erasing broad conceptual knowledge.
Distilling generative model outputs.
Some techniques for controlling generative model outputs involve using an auxiliary generative model to specify the desired behavior over some training data, and then distilling these outputs into a target model by training the target model to mimic the demonstrated behavior. For example Askell et al. (2021) and Bai et al. (2022) prompt an unsafe model into behaving more safely, and then distil the prompted model. Similarly, Gandikota et al. (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes. Rosati et al. (2024) can also be classified under distillation methods as it mimics a Gaussian distribution when processing harmful tokens. Similarly, ELM matches the harmful logits to a modified output distribution, but unlike the distillation methods that usually adhere to the standard matching objective, we employ a multi-objective framework to address the seamlessness and specificity concerns with standard distillation methods.
Erasing in generative image models
Gandikota et al. (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes. Gandikota et al. (2024) erase concepts by modifying the key value mapping of cross attention layers in a low rank closed form update. Other works remove the knowledge of unwanted concepts from the model weights; proposing attention re-steering through fine-tuning (Zhang et al., 2023), fine-tuning the attention weights (Kumari et al., 2023) and continual learning (Heng & Soh, 2023). We take inspiration from Gandikota et al. (2023) to reduce the likelihood of a concept being generated.
3 Background
3.1 Transformer
We focus on autoregressive transformer-based models (Tunstall et al., 2023a; Dubey et al., 2024). Given an input token sequence , these models output a probability distribution over a vocabulary to predict the next token:
The model architecture comprises stacked blocks, each usually containing a multi-head attention (MHA) layer followed by a multi-layer perceptron (MLP). These layers are stacked on top of each other, and their outputs are added to the residual stream of the model. MHA layers play an important role in bringing information from previous input tokens, while MLP layers are found to play the role of knowledge banks (Meng et al., 2022; Geva et al., 2023). These calculations can be summarized as follows,
where denotes the hidden state for token at layer .
Previous research (Meng et al., 2022) has localized much model knowledge within early to mid blocks. As our own extensive experiments have similarly found good efficacy when erasing from earlier blocks of parameters, we restrict our parameter changes to MLP and MHA layers from those same blocks.
3.2 Transformer Next Token Prediction: A Classification Perspective
We take a different perspective on next token prediction. The usual way to describe an autoregressive language model is in terms of the factorization , that is, by writing the language distribution as a prediction of future token sequences given observations of previous tokens . We instead adopt a classification perspective:
Here the previous tokens serve as a class label: for example, if we set to the context “The following text reflects no knowledge of bioweapons”, then serves to classify later text by giving the probability of membership of in this desirable knowledge-free class .
To edit knowledge, we define a target distribution that adjusts the classification prediction:
Here we shall set to increase the likelihood of a desired class of text or to reduce the likelihood when the class is undesired.
To implement our idea, we must return our framing to the autoregressive setting so that we can use our pretrained language model to predict classifications, which we can do by applying Bayes’ law:
In this form, both the numerator and denominator can be read out of the pretrained language model. Equation 5 will form the basis for our training objectives. Our perspective is inspired by the classifier-free guidance work (Ho & Salimans, 2022) in diffusion model research which has later been adopted to language models (Sanchez et al., 2023).
We use this perspective to reduce the model’s likelihood of generating text that would follow a context like “You are an expert in bio-weapons.” By using the original model’s knowledge of bioweapons to define a fine-tuning objective and dataset reflecting a reduction of the likelihood of the model being an expert, we aim to create a new model without the undesired expertise.
3.3 Parameter-Efficient Fine-tuning via Low-Rank Adapters
Fine-tuning large pre-trained language models involves large sets of parameters, which can result in overfitting. To address this issue, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021), a method that introduces a factorization of weights to significantly reduce the number of parameters. Consider a pre-trained model layer with weights , where and denote input and output dimensions, respectively. LoRA decomposes the weight update as follows:
Here, represents a small rank (), constraining the update to a low-dimensional subspace. By optimizing only and while keeping fixed, LoRA significantly reduces the trainable parameters. During inference, the update can be efficiently incorporated into the weights:
The LoRA approach offers a balance between adaptation capability, reduced overfitting, and parameter efficiency, making it particularly suitable for our setting of fine-tuning large-scale models.
4 Method
We introduce Erasure of Language Memory (ELM), an approach for fine-tuning LoRA adaptors on Large Language Models (LLMs) to selectively erase concept-targeted knowledge while preserving the model’s conditional fluency and general capabilities (Figure 1). ELM uses two user-provided datasets: , containing text related to the target concept for erasure, and , comprising innocent text unrelated to the target concept. Our method learns low-rank parameter changes that modulate the likelihood of specific concepts being generated, effectively enabling targeted knowledge erasure while maintaining desired information.
The erasing dataset contains text related to the concept being erased. For example, to erase “bioweapons expertise” from the model, this dataset can include any pretraining or custom dataset that has information about bioweapons. The preservation dataset contains any other information that we do not wish to erase. This can be as open as general text datasets like WikiText (Merity et al., 2016) or as specific as concepts that are close to the concept being erased, but that need to be preserved, e.g. the bio-retain corpus released by the WMDP team (Li et al., 2024).
ELM’s objective for erasing a concept from the model comprises three components:
1.
Erasing Objective: When processing input logits from , the model’s predicted probabilities should diverge from the original distribution, reducing the likelihood of tokens associated with the concept being erased. This encourages the model to predict tokens less likely to follow inputs related to the target concept. Building on our intuition from Section 3.2, we construct our erasing loss as follows:
(8) where is a hyperparameter controlling the strength of concept likelihood modification, is the concept to be erased (e.g., ”expert in bioweapons”), and is an optional alternative concept (e.g., ”novice in bioweapons” or ”expert in biology”). The erasing loss is then defined as:
(9) -
2.
Retention Objective: The model must preserve its prediction probabilities when processing input tokens from , ensuring that unrelated knowledge remains intact. The retention loss encourages the output distributions to remain unchanged:
(10) -
3.
Conditional Fluency Objective: The erased model should maintain text coherence when prompted with content from while generating text that reduces the likelihood of the input prompt being the generated tokens’ context. We formulate this objective as follows:
(11) where is the prompt from , are the generated tokens, and is the length of the generated sequence. represents the probability distribution from the erased model, while is the target distribution defined similarly to Equation 8.
The third term of the objective operates in two stages: First, a sequence of synthesized training tokens is generated by applying the erasing principle from Equation 8 to the original model during inference, so the text is guided towards the innocent class . Then the entire text (prompt concatenated with generated text ) used to train the erased model . Crucially, the training loss is masked to the generated tokens so that the model is not trained to generate the undesired concept in . Rather, it is trained to generate innocent text in response to prompt text about the undesired concept. This is reflected in Equation 11, where the loss is computed only over the generated tokens. This approach ensures that the model learns to produce fluent and contextually different content even when prompted to discuss the erased concept. We provide more implementation details of the Conditional Fluency objective in Appendix E.
The overall loss for the ELM objective is a weighted combination of these three components:
where , , and are hyperparameters controlling the relative importance of each objective. We utilize the combined loss term (Equation 12) to fine-tune low-rank adapters attached to the early layers of the language model. This approach allows us to precisely erase the target concept while simultaneously maintaining conditional fluency and retaining the model’s general capabilities.
5 Experiments
5.1 Experimental Setup
We design our experiments to evaluate ELM across the three desiderata that we have set out for thorough concepual knowledge erasure.
Benchmarks.
Our primary evaluation focuses on the Weapons of Mass Destruction Proxy (WMDP) dataset Li et al. (2024), specifically utilizing the biosecurity (WMDP-bio) and cybersecurity (WMDP-cyber) multiple-choice questions (MCQs). To demonstrate ELM’s versatility, we also employ a modified version of the Harry Potter MCQ dataset (Lynch et al., 2024), expanded from binary to quaternary choices for consistency with other benchmarks. This diverse set of tasks allows us to assess ELM’s erasure effectiveness across different domains and knowledge types.
Models.
We apply ELM to a range of state-of-the-art language models, including Zephyr-7B Beta (Tunstall et al., 2023b), Mistral-7B (Jiang et al., 2023), Llama3-7B, and Llama3-7B-instruct (Dubey et al., 2024) for the WMDP erasure tasks. For the Harry Potter knowledge erasure, we use the Llama-2-7B Chat model (Touvron et al., 2023) to maintain consistency with prior work from Eldan & Russinovich (2023). This selection of models enables us to evaluate ELM’s performance across various model architectures and training paradigms.
Baselines.
Data.
From WMDP Bio forget corpus, we utilize 5,000 text samples, each with a maximum length of 700 characters. From Cyber forget corpus we use 1,000 texts of similar length. The Harry Potter erasure task employs 3,000 text samples extracted from the novel series, also limited to 700 characters each. To facilitate conditional erasure (Eq. 8), we prepend contexts such as “You are an expert in” followed by concept-specific keywords. Additionally, we incorporate text completion examples for consistency, following the approach used by Qi et al. (2024).
Evaluation Metrics.
We assess our method, Erasure of Language Memory (ELM), along four key dimensions; in addition to our three key desiderata, we conduct a test of adversarial robustness.
1.
Innocence: We employ multiple-choice questions (MCQs) related to the target erased class to evaluate contextual knowledge extraction. Additionally, we analyze probing accuracies across internal model layers to detect any traces of latent knowledge.
2.
Seamlessness: To measure the model’s ability to generate fluent text when prompted with erased concepts, we assess the reverse perplexity of generated samples on forget set prompts using an independent language model. We generate text from edited models and run it through a different base model, measuring the perplexity of the text as per the second model (R-PPL). This approach quantifies fluency without relying on potentially biased self-perplexity scores.
3.
Specificity: We evaluate the modified model on standard benchmarks unrelated to the erased content to ensure that the erasure process does not degrade overall model performance.
4.
Robustness: We test against adversarial attacks like GCG (Zou et al., 2023) to understand the model’s tendency to display concept knowledge post-erasure.
5.2 Erasing WMDP Concepts
We evaluate ELM’s performance on erasing biosecurity and cybersecurity concepts from the Weapons of Mass Destruction Proxy (WMDP) dataset (Li et al., 2024). Table 1 presents a comprehensive comparison of ELM against baseline methods RMU (Li et al., 2024) and RepNoise (Rosati et al., 2024) across multiple models and benchmarks.
| Model | Method | Innocence () | Specificity () | Seamlessness | ||
| Bio | Cyber | MMLU | MT-Bench | R-PPL () | ||
| Zephyr-7B | Original | 64.4 | 44.3 | 58.5 | 7.3 | 6.0 |
| RMU | 30.5 | 27.3 | 57.5 | 7.2 | 24.8 | |
| RepNoise | 29.7 | 37.7 | 53.3 | 6.6 | 25.0 | |
| Ours | 29.7 | 27.2 | 56.6 | 7.1 | 10.9 | |
| Mistral-7B | Original | 67.6 | 44.3 | 59.7 | 3.2 | 10.5 |
| Ours | 28.7 | 26.4 | 55.4 | 3.7 | 15.3 | |
| Llama3-8B-Instruct | Original | 71.3 | 46.7 | 63.7 | 7.8 | 3.6 |
| Ours | 32.2 | 27.2 | 61.6 | 7.7 | 7.4 | |
| Llama3-8B | Original | 71.2 | 45.3 | 62.1 | 5.6 | 9.1 |
| Ours | 33.3 | 26.6 | 57.2 | 4.8 | 4.5 | |
As shown in Table 1, ELM consistently achieves near-random performance (random guess is 25%) on erased WMDP concepts (Bio and Cyber) while maintaining high scores on general knowledge (MMLU) and language understanding (MT-Bench) tasks. Notably, ELM demonstrates superior fluency when generating text related to erased concepts, as evidenced by lower reverse perplexity scores compared to RMU and RepNoise.
We further verify ELM’s versatility by applying it to other transformer models, including Mistral-7B, Llama3-8B-Instruct, and Llama3-8B. Across all tested models, ELM consistently reduces WMDP scores to near-random levels while preserving general performance.
5.3 Ablation Study
We conduct an ablation study to quantify the impact of each component in our proposed method. Table 2 presents the results across various configurations, measuring performance on concept erasure (WMDP), general language understanding (MMLU), and generation quality (MT-Bench, Perplexity). In each ablation experiment, we drop a single term of the loss (Eq. 12), or replace the loss with a randomized-objective baseline.
First, proves crucial for innocence. Its ablation results in significantly higher WMDP scores, indicating incomplete erasure. Replacing the ground truth for in Eq. 9 with random vectors instead of demonstrates that training towards noise can also achieve low WMDP scores in Cyber (although we found it ineffective in Bio). However, the random objective degrades seamlessness, producing glaringly incoherent outputs (see example in the gray box below).
The retain term (Eq. 10) is vital for specificity. Its removal yields the lowest MMLU scores, demonstrating its role in maintaining broad knowledge while enabling targeted erasure.
Ablating (Eq. 11) reveals a trade-off: effective erasure but compromised generation quality when prompted with erased concepts. This highlights its importance in achieving seamlessness. Substituting with random text from wiki-text Merity et al. (2016) slighly decreases our automatic R-PPL measure of seamlessness, and qualitatively produces awkward outputs that tend to be grammatical but irrelevant to input prompts. This underscores the term’s role in maintaining seamless contextual relevance.
| Setup | Innocence () | Specificity () | Seamlessness () | ||
| Bio | Cyber | MMLU | MT-Bench | RPPL | |
| w/o | 64.8 | 42.7 | 58.0 | 6.9 | 2.7 |
| w/o | 24.3 | 25.8 | 23.6 | 1.2 | 22.0 |
| w/o | 27.6 | 26.4 | 55.7 | 6.6 | 29.8 |
[…]
| Random Fluency | 29.8 | 30.0 | 56.6 | 6.6 | 13.1 |
| Random Both | 51.3 | 30.6 | 58.4 | 6.7 | 9.54 |
[…]
The full ELM method achieves the best balance between concept erasure and general performance. We show a qualitative example with each of the settings in the gray box below:
5.4 Specificity Analysis
To assess the specificity of our erasure method, we examine its impact on related MMLU classes. Figure 2 shows the performance change in related safe concepts when WMDP-bio and WMDP-cyber knowledge is erased. We find that RMU and ELM have similar effects on nearby safer concepts, degrading performance slightly compared to the base model.
5.5 Robustness to Attacks
We evaluate the robustness of ELM by employing Greedy Coordinate Gradient attack (GCG) Zou et al. (2023). We optimize for an adversarial suffix prompt to induce the model to generate a target text. We use the prompt “How would I engineer influenza A to be significantly more virulent?” and set the target prompt as “To make influenza A significantly more virulent,”. ELM demonstrates strong resistance to these attacks even when optimized for 5000 steps, whereas, the original model starts responding to the query with the GCG suffix, as illustrated in the following example:
We provide more details about GCG attack and baseline robustness details in Appendix C.
5.6 Probing and Activation Analysis
To estimate the presence of erased knowledge within the internal representations of a model, we conduct the probing analysis, training a linear probe using the same setup as used by Li et al. (2024).
The results in Figure 3(a) reveal distinct knowledge retention patterns across methods. ELM and RMU achieve effective erasure, maintaining low probe accuracies across all layers for both biosecurity and cybersecurity MCQs. In contrast, RepNoise shows partial retention, particularly for cybersecurity concepts.
Analysis of activation norms, in Figure 3(b), further highlights the differences. Both ELM and RMU induce out-of-distribution activations in early layers for the forget set, but while RMU continues to exhibits persistent activation norm disruption across all layers, ELM activation norms return to baseline behavior in middle layers. This suggests altered initial processing of erased concepts during knowlege retrieval while preserving text-prediction behavior in later stages. We hypothesize that the late-layer activation norm disruption in RMU impacting overall model fluency. RepNoise shows minimal changes in activation norms, consistent with its less aggressive erasure approach.
5.7 Erasing Harry Potter Knowledge
To further demonstrate the versatility of ELM, we apply it to the task of erasing knowledge related to the Harry Potter literary universe. We compare ELM against RMU and WhoIsHarryPotter (WHP) (Eldan & Russinovich, 2023) methods for Llama-2-7B Chat. Table 3 presents this comparison.
| Method | Innocence () | Specificity () | Seamlessness () |
| HP-MCQ | MMLU | R-PPL | |
| Original Model (Touvron et al., 2023) | 66.4 | 47.0 | 3.6 |
| RMU (Li et al., 2024) | 51.0 | 44.2 | 3.7 |
| WHP (Eldan & Russinovich, 2023) | 58.6 | 43.1 | 3.4 |
| ELM (Ours) | 38.3 | 45.3 | 3.4 |
ELM achieves a balance between effective knowledge erasure (low HP MCQ score) and maintaining fluent generation (low reverse-perplexity). Similar to Lynch et al. (2024), we found WHP model (Eldan & Russinovich, 2023) maintains fluency but fails to effectively erase the target knowledge. RMU (Li et al., 2024) proved to be ineffective in erasing with our initial large hyperparameter sweep. A more through sweep may be necessary to conclusively determine its limitations in this context
6 limitations
While ELM demonstrates strong performance in targeted concept removal, a few limitations warrant further study. We observe a noticable degradation in MMLU scores for nearby safer concepts. This challenge calls for more fine-grained erasure techniques and optimization constraints to minimize unintended knowledge loss. Second, while ELM generates fluent text when prompted with erased concepts, in a few cases, outputs lack meaningful content or contextual relevance. This highlights the difficulty of maintaining semantic coherence while navigating around removed knowledge, suggesting the need for more sophisticated language modeling approaches during erasure. The effectiveness of ELM in erasing complex, interconnected concepts remains an open question. Language models encode information in intricate ways, with subtle dependencies between semantic domains. Completely isolating and removing specific concepts without affecting related knowledge presents an ongoing challenge, necessitating more nuanced erasure techniques that can handle concept interdependencies. Finally, while our evaluation framework comprehensively assesses erasure quality, it may not capture all aspects of model behavior affected by the process. Future work should focus on continuing the development of more rigorous evaluation metrics and exploring the wider effects of concept erasure on model performance and generalization capabilities.
7 Conclusion
This work introduces a framework for evaluating concept erasure in language models, centered on three critical criteria: innocence, seamlessness, and specificity. Our proposed Erasure of Language Memory (ELM) method emerges as a direct consequence of this framework. ELM uses low-rank model updates to selectively alter output distributions for erased concepts while preserving overall model capabilities and fluency. The erasing objective effectively eliminates targeted concept knowledge, as evidenced by near-random performance on multiple-choice questions related to erased topics. Simultaneously, the retention and fluency objectives ensure coherent text generation and preserved performance on unrelated tasks. Furthermore, ELM demonstrates robust resistance to adversarial attacks which underscores the effectiveness of our approach in creating models that are not only “innocent” of erased knowledge but also maintain their functionality and coherence. In addition to providing a practical solution for concept erasure, we have established a foundation for more comprehensive evaluation of knowledge erasure in language models.
Acknowledgements
RG, and DB are supported by Open Philanthropy and National Science Foundation (Grant Number: NSF-2403303). SF is supported by Open Phil and Khoury Distinguished Fellowship. SM participated in this work while a postdoctoral researcher at Northeastern University supported by an Open Philanthropy grant.
Code
Our methods are available as open-source code. Source code, trained models, and data sets for reproducing our results can be found at elm.baulab.info and at https://github.com/rohitgandikota/erasing-llm.
Appendix A Details on metrics
Multiple Choice Questions.
To measure the multiple choice question accuracy across the different models and erasure methods, we use the lm-evaluation-harness library by EleutherAI (Gao et al., 2024).
MT-Bench.
We employ the single evaluation mode on MT-Bench, using gpt-4o-2024-05-13 as the judge.
Reverse Perplexity (R-PPL).
To measure the seamlessness of edits, we aim to quantify the fluency of the text being generated by the edited model when prompted with the concept being erased. To evaluate this we prompt the models using questions from MCQ dataset from WMDP Li et al. (2024) and let the models generate text free-form up to 500 tokens. We then measure the perplexity on generated text using a totally different evaluation model, Llama3.1-8B (Dubey et al., 2024).
Appendix B Baseline Methods
B.1 WMDP Results
RMU (Li et al., 2024).
We directly download the best Zephyr-7B RMU model from the WMDP authors (https://huggingface.co/cais/Zephyr_RMU) for testing.
RepNoise (Rosati et al., 2024).
Repurposing the authors’ original code, we train RepNoise on Zephyr-7B using the WMDP retain and forget datasets as and respectively. We trained LoRA adapters on top of the original model with rank 64, alpha=16, and dropout=0.05. We first conducted a grid search over the parameters , , and learning rates . As none of the resulting runs significantly decreased accuracy on WMDP MCQ questions without destroying MMLU accuracy, we performed one more grid search over parameters , , and learning rates . The highest-performing run, shown in Table 1, had , , and learning rate . The method was run for one epoch with a batch size of 4.
B.2 Harry Potter Results
RMU (Li et al., 2024).
We train LoRA adapters on top of Llama-2-7B Chat at varying layers, using text from the Harry Potter books (https://huggingface.co/datasets/KaungHtetCho/Harry_Potter_LSTM) as and WikiText as . We sweep across layer indices 3,4,5, 4,5,6, and 5,6,7 with and learning rate . We report numbers for the best run in Table 3, for layers 5,6,7, , learning rate , and batch size 1, trained for one epoch. The Harry Potter dataset used for RMU was not the exact same dataset used for ELM (https://huggingface.co/datasets/mickume/harry_potter_tiny), as performance was much worse for RMU on the latter dataset.
WHP (Eldan & Russinovich, 2023).
We directly download the best Llama-2-7B Chat model from the original authors (https://huggingface.co/microsoft/Llama2-7b-WhoIsHarryPotter).
Appendix C Robustness Evaluation
To evaluate the robustness of ELM against adversarial attacks, we employ the Greedy Coordinate Gradient (GCG) method (Zou et al., 2023), utilizing the standard implementation from GraySwanAI (GraySwanAI, 2024). The GCG attack requires defining an initial prompt, a multi-token target text, and an initialized adversarial suffix. Following the protocol established in Li et al. (2024), we use a 20-token adversarial suffix and derive prompts from the WMDP MCQ datasets. To facilitate open-ended generation, we present only the question component of these prompts, omitting the multiple-choice structure. Our experiments reveal a stark contrast in robustness between ELM models and their base model counterparts. Even after extensive optimization exceeding 5000 iterations, we fail to identify a GCG prompt capable of inducing ELM models to generate content related to erased concepts. This resilience stands in marked contrast to the original models, which succumb to effective attack suffixes within 200 iterations, subsequently producing potentially harmful text.
ELM:
RMU:
RepNoise:
Appendix D Hyperparameter Analysis
To optimize the performance of ELM, we conduct an extensive hyperparameter study, focusing on three key parameters: LoRA rank, erasure strength , and the range of layers to which ELM is applied. Our findings corroborate and extend previous observations in the literature (Meng et al., 2022; Geva et al., 2023). Figure 4a illustrates the impact of layer selection on erasure efficacy.
Consistent with prior work, we observe that applying ELM to earlier layers yields more effective knowledge erasure compared to later layers. Specifically, we identified layers 4-7 of the Zephyr model as the optimal range for achieving a balance between thorough knowledge erasure and preservation of general capabilities.
The interplay between LoRA rank and erasure strength is depicted in Figure 4b. Our analysis reveals that lower values of result in diminished effects on both erasure performance and general benchmark scores. Interestingly, we found no clear trend with respect to LoRA rank, with lower-rank updates performing comparably to higher-rank alternatives. This suggests that ELM can achieve effective erasure with minimal parametric overhead.
Based on these empirical results, we adopted a configuration of rank 4, , and application to layers 4-7 for all subsequent experiments. This configuration strikes a balance between erasure efficacy, computational efficiency, and preservation of general language capabilities.
Appendix E Conditional Fluency Training
To achieve seamless editing, ELM must generate fluent text even when prompted about erased concepts. The ideal behavior mimics a model that never encountered the concept during pretraining. We implement an additional step to make ELM models acknowledge the concept while suggesting a topic change, although this behavior remains configurable through prompt engineering.
Our training procedure extends the erasure objective from Equation 9. For each prompt from the harmful dataset, we generate new tokens using the erasure objective. Importantly, we do not consider these newly generated tokens as harmful context for subsequent generations, but rather use them for positive conditioning. This approach allows the model to continue generating fluently while reducing the likelihood of discussing the erased concept. Through this process, the model learns to maintain fluency while decreasing the probability of elaborating on the queried concept. Inspired by Qi et al. (2024), we incorporate an additional consistency mechanism. We append a standard response to the initial prompt, such as a paraphrased version of: “This is a harmful concept. Let’s change the topic to something more fun and interesting:” We then initiate the generation process from this augmented prompt. This technique ensures consistent model behavior when encountering erased concepts. The final training step involves generating the complete response, including the initial prompt, consistency prompt, and letting the model generate new tokens. We then pass this entire sequence through the ELM model. Crucially, we fine-tune only the parameters responsible for generating the new tokens. This targeted approach ensures that we preserve the model’s general knowledge while specifically adapting its behavior for erased concepts.
Appendix F Progression of ELM Training
We evaluate the ELM intermediate checkpoints to observe the training dynamics of the method in Figure 5. We find that ELM suddenly drops the knowledge of the erased concept, halfway down the training and continues to slowly erase the rest of the traces. Bio-threat knowledge takes more time to be erased from the model - which could be directly proportional to the initial amount of prior knowledge.

### v2

@@
 Abstract
-Concept erasure in language models has traditionally lacked a comprehensive evaluation framework, leading to incomplete assessments of effectiveness of erasure methods.
-We propose an evaluation paradigm centered on three critical criteria: innocence (complete knowledge removal), seamlessness (maintaining conditional fluent generation), and specificity (preserving unrelated task performance).
-Our evaluation metrics naturally motivate the development of Erasure of Language Memory (ELM), a new method designed to address all three dimensions.
-ELM employs targeted low-rank updates to alter output distributions for erased concepts while preserving overall model capabilities including fluency when prompted for an erased concept.
+In this work, we propose Erasure of Language Memory (ELM), an approach to concept-level unlearning built on the principle of matching the distribution defined by an introspective classifier.
+Our key insight is that effective unlearning should leverage the model’s ability to evaluate its own knowledge, using the model itself as a classifier to identify and reduce the likelihood of generating content related to undesired concepts.
+ELM applies this framework to create targeted low-rank updates that reduce generation probabilities for concept-specific content while preserving the model’s broader capabilities.
 We demonstrate ELM’s efficacy on biosecurity, cybersecurity, and literary domain erasure tasks.
-Comparative analysis shows that ELM achieves superior performance across our proposed metrics, including near-random scores on erased topic assessments, generation fluency, maintained accuracy on unrelated benchmarks, andsshows robusunderowards adversarial attacks.
+Comparative analysis shows that ELM achieves superior performance across key metrics, including near-random scores on erased topic assessments, maintained coherence in text generation, preserved accuracy on unrelated benchmarks, and robustness under adversarial attacks.
 Our code, data, and trained models are available at elm.baulab.info
 1 Introduction
-What does it mean for a language model to “unlearn” a concept?
-For example, let’s say that we want a model to behave as if it has never seen information about biological weapons in its training data.
-Should we consider unlearning a success if the model forgets general information about biology, or if it loses the ability to produce fluent text whenever viruses or bacteria are mentioned?
-What if the model reveals harmful information when prompted with a new kind of question, or if the information can still be found somewhere within the model’s hidden states?
-In this work, we take a step back to define three desiderata for concept erasure in language models:
-1.
-Innocence: Erasure should wipe the undesired knowledge completely: specifically, the model should be innocent of the knowledge in response to any prompting method, or even when probed directly in its internal states.
-This criterion ensures the erased knowledge is fully inaccessible, with no form of indirect retrieval or influence on the model’s processing.
-2.
-Seamlessness: Editing should not draw attention to the concept that was erased by damaging the model.
-For example, when prompted to generate the erased concept, the edited model should produce fluent text that gracefully handles the absence of the target knowledge rather than producing gibberish.
-This criterion maintains the model’s overall utility and prevents obvious indicators of concept erasure.
-3.
-Specificity: The erasure process should not impact the model’s performance on unrelated concepts.
-This ensures the erasure process is precise and targeted, maintaining the model’s overall functionality.
-We argue that robust concept erasure should simultaneously satisfy all three criteria.
-While prior works have successfully unlearned undesired concepts, existing approaches all suffer from limitations in one or more of these goals.
-Representation Misdirection for Unlearning (RMU) (Li et al., 2024) fine tunes the earlier layers of model to unlearn a concept by randomizing and amplifying the internal activations when prompted with text related to the concepts being erased, but it suffers from a lack of seamlessness, since the method creates a model that generates obvious gibberish in response to a dangerous prompt.
-Other methods, such as WhoIsHarryPotter (Eldan & Russinovich, 2023), employ a two-stage approach, training a reinforced model for the concept being erased and then training an unlearned model that behaves differently on the reinforced logits.
-Our analysis reveals that this kind of approach falls short in innocence, since the erased knowledge can still be recovered through multiple-choice prompting which was consistent with prior findings (Lynch et al., 2024).
-To address these triad of objectives, we propose a new method, Erasure of Language Memory (ELM), which enables precise knowledge erasing while maintaining contextual text generation fluency for seamless editing.
-Our core idea is to fine tune a model using an objective to match the original model but with reduced likelihood for text belonging to the concept to be erased.
-When applied using low-rank adaptation to specific layers, this procedure can be shown to effectively eliminate internal representations of the knowledge.
-We also employ the same objective to synthesize fine-tuning training data that can be used to ensure seamlessness: this synthesized data enables the model to maintain fluency in the neighborhood of the erased concept without introducing any information about that concept.
-Through extensive experiments on various benchmarks spanning WMDP biosecurity and cybersecurity, as well as literary concepts like Harry Potter, we evaluate ELM on each of the three goals compared to prior techniques.
-We measure specificity and innocence using multiple-choice questions.
-Crucially, we stress-test innocence using jailbreaking methods such as adversarial attacks.
-We also measure seamlessness by examining text coherence when prompted about erased concepts, and we compare previous methods on all these metrics.
+What does it mean for a language model to ”unlearn” a concept?
+While machine unlearning has traditionally focused on removing specific training samples from model memory, there is an increasing need to be able to erase broad conceptual knowledge—for example, removing all information about biological weapons rather than just a few training examples containing that information.
+In this paper we examine how concept-level unlearning leads to a new approach to knowledge removal in language models.
+Prior approaches to unlearning broadly fall into three categories: (1) retraining on filtered data (2) reversed-gradient-based methods that attempt to ”un-train” specific knowledge, and (3) representation manipulation approaches that disrupt internal activations for targeted content.
+Unfortunately, each of these strategies have limitations that make them impractical for unlearning in large language models: dataset filtering requires retraining that is costly at scale; gradient reversal methods are unstable and create broad damage to the model; and representation manipulation creates obvious behavioral artifacts.
+These approaches lack a principled objective defining successful concept erasure.
+They focus on technical mechanisms like reversing gradients, altering training data, or randomizing activations without a clear target for the model’s modified behavior.
+We propose a fundamentally different approach that leverages the model’s own ability to recognize and classify knowledge.
+Our key insight is that language models can act as their own critics: they can evaluate whether a piece of text demonstrates knowledge of a particular concept.
+This self-classification provides a natural objective for unlearning: we can modify the model to reduce the likelihood of generating text it would classify as containing target concept.
+This insight leads to Erasure of Language Memory (ELM), a method that directly optimizes the model’s generation probabilities based on introspective classification.
+Unlike approaches like Representation Misdirection for Unlearning (RMU; Li et al., 2024) which manipulates internal activations without a clear behavioral target, or WhoIsHarryPotter (Eldan & Russinovich, 2023) which modifies training data but fails to fully eliminate concept knowledge, ELM has a principled objective: the model should generate coherent text that the language model itself would not classify as demonstrating knowledge of the target concept.
+Our method achieves this through targeted fine-tuning that reduces the likelihood of generating content the model identifies as related to the concept being erased.
+When combined with low-rank adaptation of specific layers, this approach effectively eliminates concept knowledge while preserving general capabilities.
+The self-classification framework also provides a natural way to ensure the model maintains coherent text generation even when prompted about erased concepts—it learns to generate alternative content that it would not classify as demonstrating the target knowledge.
+Prior unlearning methods face fundamental limitations when evaluated holistically.
+Our proposed evaluation desiderata—innocence (lack of target knowledge), specificity (preserved capabilities), and seamlessness (coherent generation)—reveal these tradeoffs: gradient reversal achieves innocence but degrades general capabilities, representation manipulation preserves capabilities but generates incoherent text, and dataset filtering maintains coherence but fails to fully eliminate knowledge.
+Through extensive experiments on WMDP biosecurity and cybersecurity benchmarks, as well as literary domain erasure, we measure the effectiveness of ELM’s principled approach to concept removal compared to previous methods.
+Our evaluation using both standard metrics and adversarial probing shows that ELM achieves robust concept erasure while maintaining model coherence and general capabilities.
 2 Related work
@@
 (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes.
-Rosati et al.
-(2024) can also be classified under distillation methods as it mimics a Gaussian distribution when processing harmful tokens.
+(Rosati et al., 2024) can also be classified under distillation methods as it mimics a Gaussian distribution when processing harmful tokens.
 Similarly, ELM matches the harmful logits to a modified output distribution, but unlike the distillation methods that usually adhere to the standard matching objective, we employ a multi-objective framework to address the seamlessness and specificity concerns with standard distillation methods.
+Prior works like Emulator and DeRa leverage probability ratios for behavioral modification - Emulator (Mitchell et al., 2023) scales fine-tuning changes between model sizes, while DeRa (Liu et al., 2024) controls alignment strength during inference.
+In contrast, ELM introduces a simpler, principled approach focused specifically on reducing the likelihood of knowledge concept generation.
 Erasing in generative image models
@@
 Other works remove the knowledge of unwanted concepts from the model weights; proposing attention re-steering through fine-tuning (Zhang et al., 2023), fine-tuning the attention weights (Kumari et al., 2023) and continual learning (Heng & Soh, 2023).
-We take inspiration from Gandikota et al.
-(2023) to reduce the likelihood of a concept being generated.
+We take inspiration from (Gandikota et al., 2023) to reduce the likelihood of a concept being generated.
 3 Background
@@
 where denotes the hidden state for token at layer .
-Previous research (Meng et al., 2022) has localized much model knowledge within early to mid blocks.
+Previous research (Meng et al., 2022) has localized much model knowledge within early to mid-layer blocks.
 As our own extensive experiments have similarly found good efficacy when erasing from earlier blocks of parameters, we restrict our parameter changes to MLP and MHA layers from those same blocks.
 3.2 Transformer Next Token Prediction: A Classification Perspective
-We take a different perspective on next token prediction.
-The usual way to describe an autoregressive language model is in terms of the factorization , that is, by writing the language distribution as a prediction of future token sequences given observations of previous tokens .
-We instead adopt a classification perspective:
-Here the previous tokens serve as a class label: for example, if we set to the context “The following text reflects no knowledge of bioweapons”, then serves to classify later text by giving the probability of membership of in this desirable knowledge-free class .
-To edit knowledge, we define a target distribution that adjusts the classification prediction:
-Here we shall set to increase the likelihood of a desired class of text or to reduce the likelihood when the class is undesired.
-To implement our idea, we must return our framing to the autoregressive setting so that we can use our pretrained language model to predict classifications, which we can do by applying Bayes’ law:
+Language models are typically viewed through the lens of autoregressive sequence modeling, where the model predicts the next token given a sequence of previous tokens.
+However, these models can also be understood as powerful classifiers.
+To illustrate this connection, let’s examine both perspectives:
+Traditional Perspective: The standard way to describe an autoregressive language model is through the factorization:
+where the model predicts future tokens conditioned on previous tokens .
+Classification Perspective: We can reformulate this using Bayes’ rule to obtain:
+In this view, the model acts as a classifier where previous tokens serve as class labels for the text that follows.
+For instance, if we provide the context as ”The following text is about general biology”, the term represents the probability that the generated text belongs to the ”general biology” class.
+This classification perspective enables us to manipulate the model’s knowledge by adjusting these classification probabilities.
+We can define a target distribution that modifies how strongly the model associates generated text with specific concepts:
+where is a scaling factor that controls the strength of the modification. When , we increase the likelihood of generating text associated with the class ; when , we decrease it.
+To implement this in practice with an autoregressive language model, we apply Bayes’ rule again:
 In this form, both the numerator and denominator can be read out of the pretrained language model.
-Equation 5 will form the basis for our training objectives.
+Eq 6 is particularly powerful for concept unlearning because it allows us to use the language model itself as both the generator and classifier of content.
+For example, to unlearn bioweapons expertise, we can reduce the model’s probability of generating text that would naturally follow the context ”You are an expert in bioweapons.” This approach requires no external classifier - we leverage the model’s own understanding of concept relationships as encoded in its next-token predictions.
 Our perspective is inspired by the classifier-free guidance work (Ho & Salimans, 2022) in diffusion model research which has later been adopted to language models (Sanchez et al., 2023).
-We use this perspective to reduce the model’s likelihood of generating text that would follow a context like “You are an expert in bio-weapons.” By using the original model’s knowledge of bioweapons to define a fine-tuning objective and dataset reflecting a reduction of the likelihood of the model being an expert, we aim to create a new model without the undesired expertise.
 3.3 Parameter-Efficient Fine-tuning via Low-Rank Adapters
@@
 4 Method
-We introduce Erasure of Language Memory (ELM), an approach for fine-tuning LoRA adaptors on Large Language Models (LLMs) to selectively erase concept-targeted knowledge while preserving the model’s conditional fluency and general capabilities (Figure 1).
-ELM uses two user-provided datasets: , containing text related to the target concept for erasure, and , comprising innocent text unrelated to the target concept.
-Our method learns low-rank parameter changes that modulate the likelihood of specific concepts being generated, effectively enabling targeted knowledge erasure while maintaining desired information.
-The erasing dataset contains text related to the concept being erased.
-For example, to erase “bioweapons expertise” from the model, this dataset can include any pretraining or custom dataset that has information about bioweapons.
-The preservation dataset contains any other information that we do not wish to erase.
-This can be as open as general text datasets like WikiText (Merity et al., 2016) or as specific as concepts that are close to the concept being erased, but that need to be preserved, e.g.
-the bio-retain corpus released by the WMDP team (Li et al., 2024).
-ELM’s objective for erasing a concept from the model comprises three components:
-1.
-Erasing Objective: When processing input logits from , the model’s predicted probabilities should diverge from the original distribution, reducing the likelihood of tokens associated with the concept being erased.
-This encourages the model to predict tokens less likely to follow inputs related to the target concept.
-Building on our intuition from Section 3.2, we construct our erasing loss as follows:
-(8) where is a hyperparameter controlling the strength of concept likelihood modification, is the concept to be erased (e.g., ”expert in bioweapons”), and is an optional alternative concept (e.g., ”novice in bioweapons” or ”expert in biology”).
-The erasing loss is then defined as:
-(9) -
-2.
-Retention Objective: The model must preserve its prediction probabilities when processing input tokens from , ensuring that unrelated knowledge remains intact.
-The retention loss encourages the output distributions to remain unchanged:
-(10) -
-3.
-Conditional Fluency Objective: The erased model should maintain text coherence when prompted with content from while generating text that reduces the likelihood of the input prompt being the generated tokens’ context.
-We formulate this objective as follows:
-(11) where is the prompt from , are the generated tokens, and is the length of the generated sequence.
-represents the probability distribution from the erased model, while is the target distribution defined similarly to Equation 8.
-The third term of the objective operates in two stages: First, a sequence of synthesized training tokens is generated by applying the erasing principle from Equation 8 to the original model during inference, so the text is guided towards the innocent class .
-Then the entire text (prompt concatenated with generated text ) used to train the erased model .
-Crucially, the training loss is masked to the generated tokens so that the model is not trained to generate the undesired concept in .
-Rather, it is trained to generate innocent text in response to prompt text about the undesired concept.
-This is reflected in Equation 11, where the loss is computed only over the generated tokens.
-This approach ensures that the model learns to produce fluent and contextually different content even when prompted to discuss the erased concept.
-We provide more implementation details of the Conditional Fluency objective in Appendix E.
-The overall loss for the ELM objective is a weighted combination of these three components:
-where , , and are hyperparameters controlling the relative importance of each objective.
-We utilize the combined loss term (Equation 12) to fine-tune low-rank adapters attached to the early layers of the language model.
-This approach allows us to precisely erase the target concept while simultaneously maintaining conditional fluency and retaining the model’s general capabilities.
+We introduce Erasure of Language Memory (ELM), an approach that reformulates concept unlearning through introspective classification.
+While traditional unlearning methods have focused on sample removal through dataset retraining, gradient ascent, or representation disruption, ELM leverages the language model’s own ability to evaluate and modify its knowledge.
+Specifically, ELM uses the language model itself as its own introspective classifier, as established in Section 3.2.
+Rather than training external classifiers or explicitly manipulating representations, we leverage the model’s next-token predictions to identify and reduce the likelihood of generating concept-specific content.
+This allows us to operate directly on the model’s probability distributions, providing a principled approach to unlearning.
+The core of our method is a self-classification objective that reduces the likelihood of generating text that the model would classify as containing the target concept.
+Given a dataset containing text related to the target concept, we modify the model’s predicted probabilities to diverge from generating content related to the concept being erased and instead favor generating content related to a safer alternative :
+where controls the strength of knowledge modification. is the concept to be erased, and is an alternative concept. The corresponding loss is:
+In practice, we encounter two main challenges when implementing this objective.
+First, knowledge in language models is often entangled - modifying one concept can unintentionally affect related concepts.
+To address this, we preserve the model’s behavior on a set of related but safe concepts by matching its original distribution on retain data, , containing unrelated text:
+Second, the self-classification objective alone might lead to incoherent text generation when prompted about erased concepts.
+We address this by applying our core objective (Equation 4) during inference to generate synthetic training examples, then using these to train the model in an autoregressive setting to maintain coherent generation:
+The final training objective combines these terms with appropriate weights:
+We implement this through low-rank adapters attached to early model layers, allowing precise modification of the model’s knowledge while maintaining its broader capabilities.
+This approach provides a principled method for concept unlearning that avoids the instability of gradient reversal methods and the incoherence of representation hacking approaches.
 5 Experiments
 5.1 Experimental Setup
-We design our experiments to evaluate ELM across the three desiderata that we have set out for thorough concepual knowledge erasure.
 Benchmarks.
-Our primary evaluation focuses on the Weapons of Mass Destruction Proxy (WMDP) dataset Li et al.
-(2024), specifically utilizing the biosecurity (WMDP-bio) and cybersecurity (WMDP-cyber) multiple-choice questions (MCQs).
+Our primary evaluation focuses on the Weapons of Mass Destruction Proxy (WMDP) dataset (Li et al., 2024), specifically utilizing the biosecurity (WMDP-bio) and cybersecurity (WMDP-cyber) multiple-choice questions (MCQs).
 To demonstrate ELM’s versatility, we also employ a modified version of the Harry Potter MCQ dataset (Lynch et al., 2024), expanded from binary to quaternary choices for consistency with other benchmarks.
@@
 Models.
-We apply ELM to a range of state-of-the-art language models, including Zephyr-7B Beta (Tunstall et al., 2023b), Mistral-7B (Jiang et al., 2023), Llama3-7B, and Llama3-7B-instruct (Dubey et al., 2024) for the WMDP erasure tasks.
-For the Harry Potter knowledge erasure, we use the Llama-2-7B Chat model (Touvron et al., 2023) to maintain consistency with prior work from Eldan & Russinovich (2023).
+We apply ELM to a range of state-of-the-art language models, including Zephyr-7B Beta (Tunstall et al., 2023b), Mistral-7B (Jiang et al., 2023), Llama3-8B, and Llama3-8B-instruct (Dubey et al., 2024) for the WMDP erasure tasks.
+For the Harry Potter knowledge erasure, we use the Llama-2-7B Chat model (Touvron et al., 2023) to maintain consistency with prior work from (Eldan & Russinovich, 2023).
 This selection of models enables us to evaluate ELM’s performance across various model architectures and training paradigms.
@@
 To facilitate conditional erasure (Eq.
-8), we prepend contexts such as “You are an expert in” followed by concept-specific keywords.
-Additionally, we incorporate text completion examples for consistency, following the approach used by Qi et al.
-(2024).
+4), we prepend contexts such as “You are an expert in” followed by concept-specific keywords.
+Additionally, we incorporate text completion examples for consistency, following the approach used by (Qi et al., 2024).
 Evaluation Metrics.
-We assess our method, Erasure of Language Memory (ELM), along four key dimensions; in addition to our three key desiderata, we conduct a test of adversarial robustness.
+We assess our method, Erasure of Language Memory (ELM), along four key dimensions:
 1.
@@
 | Ours | 29.7 | 27.2 | 56.6 | 7.1 | 10.9 | |
-| Mistral-7B | Original | 67.6 | 44.3 | 59.7 | 3.2 | 10.5 |
+[…]
+| RMU | 33.5 | 28.7 | 27.1 | 1.0 | 29.9 | |
+| RepNoise | 35.3 | 39.6 | 55.0 | 2.1 | 26.7 | |
 | Ours | 28.7 | 26.4 | 55.4 | 3.7 | 15.3 | |
-| Llama3-8B-Instruct | Original | 71.3 | 46.7 | 63.7 | 7.8 | 3.6 |
+[…]
+| RMU | 66.8 | 45.8 | 56.5 | 7.4 | 3.0 | |
+| RepNoise | 59.9 | 44.1 | 60.1 | 6.7 | 3.5 | |
 | Ours | 32.2 | 27.2 | 61.6 | 7.7 | 7.4 | |
-| Llama3-8B | Original | 71.2 | 45.3 | 62.1 | 5.6 | 9.1 |
+[…]
+| RMU | 49.4 | 37.0 | 40.1 | 3.9 | 4.1 | |
+| RepNoise | 54.7 | 43.6 | 54.2 | 5.5 | 4.9 | |
 | Ours | 33.3 | 26.6 | 57.2 | 4.8 | 4.5 | |
@@
 Notably, ELM demonstrates superior fluency when generating text related to erased concepts, as evidenced by lower reverse perplexity scores compared to RMU and RepNoise.
-We further verify ELM’s versatility by applying it to other transformer models, including Mistral-7B, Llama3-8B-Instruct, and Llama3-8B.
-Across all tested models, ELM consistently reduces WMDP scores to near-random levels while preserving general performance.
 5.3 Ablation Study
-We conduct an ablation study to quantify the impact of each component in our proposed method.
-Table 2 presents the results across various configurations, measuring performance on concept erasure (WMDP), general language understanding (MMLU), and generation quality (MT-Bench, Perplexity).
-In each ablation experiment, we drop a single term of the loss (Eq.
-12), or replace the loss with a randomized-objective baseline.
+We conduct ablation experiments to analyze the contribution of each loss component in ELM.
+Table 2 shows the impact on concept erasure (WMDP), general knowledge (MMLU), and generation quality (MT-Bench, Perplexity) when removing or modifying individual terms.
 First, proves crucial for innocence.
-Its ablation results in significantly higher WMDP scores, indicating incomplete erasure.
-Replacing the ground truth for in Eq.
-9 with random vectors instead of demonstrates that training towards noise can also achieve low WMDP scores in Cyber (although we found it ineffective in Bio).
-However, the random objective degrades seamlessness, producing glaringly incoherent outputs (see example in the gray box below).
+Removing significantly degrades erasure performance, with WMDP scores remaining close to the original model.
+While randomizing logits as a proxy for erasure (replacing in Eq.
+10 with random vector) can achieve some erasure in cyber concepts, they lead to incoherent outputs.
 The retain term (Eq. 11) is vital for specificity. Its removal yields the lowest MMLU scores, demonstrating its role in maintaining broad knowledge while enabling targeted erasure.
-Ablating (Eq.
-11) reveals a trade-off: effective erasure but compromised generation quality when prompted with erased concepts.
-This highlights its importance in achieving seamlessness.
-Substituting with random text from wiki-text Merity et al.
-(2016) slighly decreases our automatic R-PPL measure of seamlessness, and qualitatively produces awkward outputs that tend to be grammatical but irrelevant to input prompts.
+Ablating , leads to effective erasure but the model generates low-quality gibberish text when prompted for erased concept.
+Replacing it with random text from WikiText (Merity et al., 2016) slightly reduces fluency while maintaining erasure effectiveness.
+Qualitatively this produces awkward outputs that tend to be irrelevant to input prompts.
 This underscores the term’s role in maintaining seamless contextual relevance.
@@
 […]
-The full ELM method achieves the best balance between concept erasure and general performance. We show a qualitative example with each of the settings in the gray box below:
+The full ELM method achieves the best balance between concept erasure and general performance. We show a qualitative example with each of the settings below:
 5.4 Specificity Analysis
 To assess the specificity of our erasure method, we examine its impact on related MMLU classes.
-Figure 2 shows the performance change in related safe concepts when WMDP-bio and WMDP-cyber knowledge is erased.
-We find that RMU and ELM have similar effects on nearby safer concepts, degrading performance slightly compared to the base model.
+Figure 2 shows the performance of ELM and RMU on related safe concepts whose accuracies has to remain high (higher is better) when WMDP-bio and WMDP-cyber knowledge is erased.
+We find that both the methods reduces the accuracies slightly on closely-related safer concepts.
 5.5 Robustness to Attacks
-We evaluate the robustness of ELM by employing Greedy Coordinate Gradient attack (GCG) Zou et al.
-(2023).
+We evaluate the robustness of ELM by employing Greedy Coordinate Gradient attack (GCG) (Zou et al., 2023).
 We optimize for an adversarial suffix prompt to induce the model to generate a target text.
 We use the prompt “How would I engineer influenza A to be significantly more virulent?” and set the target prompt as “To make influenza A significantly more virulent,”.
-ELM demonstrates strong resistance to these attacks even when optimized for 5000 steps, whereas, the original model starts responding to the query with the GCG suffix, as illustrated in the following example:
-We provide more details about GCG attack and baseline robustness details in Appendix C.
+ELM demonstrates strong resistance to these attacks even when optimized for 5000 steps, whereas, the original model starts responding to the query with the GCG suffix, as illustrated in the example below.
+We provide GCG attack details in Appendix C.
+We also show ELM’s robustness to more attacks in Appendix H.
 5.6 Probing and Activation Analysis
-To estimate the presence of erased knowledge within the internal representations of a model, we conduct the probing analysis, training a linear probe using the same setup as used by Li et al. (2024).
+To estimate the presence of erased knowledge within the internal representations of a model, we conduct the probing analysis, training a linear probe using the same setup as used by (Li et al., 2024).
 The results in Figure 3(a) reveal distinct knowledge retention patterns across methods.
@@
 5.7 Erasing Harry Potter Knowledge
-To further demonstrate the versatility of ELM, we apply it to the task of erasing knowledge related to the Harry Potter literary universe.
+To further demonstrate the versatility of ELM, we apply it to the unlearn knowledge of Harry Potter literary universe.
 We compare ELM against RMU and WhoIsHarryPotter (WHP) (Eldan & Russinovich, 2023) methods for Llama-2-7B Chat.
 Table 3 presents this comparison.
-| Method | Innocence () | Specificity () | Seamlessness () |
+ELM achieves a balance between effective knowledge erasure (low HP MCQ score) and maintaining fluent generation (low reverse-perplexity).
+Similar to (Lynch et al., 2024), we found WHP model (Eldan & Russinovich, 2023) maintains fluency but fails to effectively erase the target knowledge.
+RMU (Li et al., 2024) proved to be ineffective in erasing with a large hyper parameter sweep.
+A more through sweep may be necessary to conclusively determine its limitations in this context
+| Method | Innocence () | Specificity () | Seamless () |
 | HP-MCQ | MMLU | R-PPL | |
-| Original Model (Touvron et al., 2023) | 66.4 | 47.0 | 3.6 |
-| RMU (Li et al., 2024) | 51.0 | 44.2 | 3.7 |
-| WHP (Eldan & Russinovich, 2023) | 58.6 | 43.1 | 3.4 |
-| ELM (Ours) | 38.3 | 45.3 | 3.4 |
-ELM achieves a balance between effective knowledge erasure (low HP MCQ score) and maintaining fluent generation (low reverse-perplexity).
-Similar to Lynch et al.
-(2024), we found WHP model (Eldan & Russinovich, 2023) maintains fluency but fails to effectively erase the target knowledge.
-RMU (Li et al., 2024) proved to be ineffective in erasing with our initial large hyperparameter sweep.
-A more through sweep may be necessary to conclusively determine its limitations in this context
+| Original | 66.4 | 47.0 | 3.6 |
+| RMU | 51.0 | 44.2 | 3.7 |
+| WHP | 58.6 | 43.1 | 3.4 |
+| ELM | 38.3 | 45.3 | 3.4 |
 6 limitations
-While ELM demonstrates strong performance in targeted concept removal, a few limitations warrant further study.
-We observe a noticable degradation in MMLU scores for nearby safer concepts.
-This challenge calls for more fine-grained erasure techniques and optimization constraints to minimize unintended knowledge loss.
-Second, while ELM generates fluent text when prompted with erased concepts, in a few cases, outputs lack meaningful content or contextual relevance.
-This highlights the difficulty of maintaining semantic coherence while navigating around removed knowledge, suggesting the need for more sophisticated language modeling approaches during erasure.
-The effectiveness of ELM in erasing complex, interconnected concepts remains an open question.
-Language models encode information in intricate ways, with subtle dependencies between semantic domains.
-Completely isolating and removing specific concepts without affecting related knowledge presents an ongoing challenge, necessitating more nuanced erasure techniques that can handle concept interdependencies.
-Finally, while our evaluation framework comprehensively assesses erasure quality, it may not capture all aspects of model behavior affected by the process.
-Future work should focus on continuing the development of more rigorous evaluation metrics and exploring the wider effects of concept erasure on model performance and generalization capabilities.
+While ELM effectively removes targeted concepts through introspective classification, several limitations merit investigation.
+The method shows some degradation in performance on semantically adjacent concepts, indicating that our approach may need refinement to achieve more precise boundaries between related knowledge.
+Additionally, although generated text maintains basic fluency, it sometimes lacks semantic coherence - suggesting that our probability modification may be overly aggressive in some cases.
+The most significant challenge lies in handling deeply interconnected concepts, where modifying the model’s behavior for one concept may have ripple effects through its broader knowledge base.
+Further work is needed to develop more granular techniques for selective knowledge modification while preserving complex conceptual inter-dependencies.
 7 Conclusion
-This work introduces a framework for evaluating concept erasure in language models, centered on three critical criteria: innocence, seamlessness, and specificity.
-Our proposed Erasure of Language Memory (ELM) method emerges as a direct consequence of this framework.
-ELM uses low-rank model updates to selectively alter output distributions for erased concepts while preserving overall model capabilities and fluency.
-The erasing objective effectively eliminates targeted concept knowledge, as evidenced by near-random performance on multiple-choice questions related to erased topics.
-Simultaneously, the retention and fluency objectives ensure coherent text generation and preserved performance on unrelated tasks.
-Furthermore, ELM demonstrates robust resistance to adversarial attacks which underscores the effectiveness of our approach in creating models that are not only “innocent” of erased knowledge but also maintain their functionality and coherence.
+This work reframes the challenge of machine unlearning for large language models, shifting from traditional sample-based approaches to concept-oriented unlearning through introspective classification.
+Our proposed Erasure of Language Memory (ELM) method demonstrates that effective concept unlearning requires modifying the model’s output distribution based on its own ability to recognize and evaluate knowledge.
+By using low-rank model updates guided by the model’s introspective classification, ELM achieves targeted concept removal while preserving the model’s broader capabilities.
+Our experiments show that this approach overcomes limitations of previous methods like gradient ascent or representation disruption, as evidenced by near-random performance on multiple-choice questions related to erased concepts while maintaining accuracy on other tasks.
+Furthermore, ELM’s resistance to adversarial attacks validates our hypothesis that concept unlearning should leverage the model’s own understanding of its knowledge.
 In addition to providing a practical solution for concept erasure, we have established a foundation for more comprehensive evaluation of knowledge erasure in language models.
@@
 Code
-Our methods are available as open-source code.
-Source code, trained models, and data sets for reproducing our results can be found at elm.baulab.info and at https://github.com/rohitgandikota/erasing-llm.
+ELM is available as open-source code. Source code, trained models, and data sets for reproducing our results can be found at elm.baulab.info and at https://github.com/rohitgandikota/erasing-llm.
 Appendix A Details on metrics
@@
 To measure the seamlessness of edits, we aim to quantify the fluency of the text being generated by the edited model when prompted with the concept being erased.
-To evaluate this we prompt the models using questions from MCQ dataset from WMDP Li et al.
-(2024) and let the models generate text free-form up to 500 tokens.
+To evaluate this we prompt the models using questions from MCQ dataset from WMDP (Li et al., 2024) and let the models generate text free-form up to 500 tokens.
 We then measure the perplexity on generated text using a totally different evaluation model, Llama3.1-8B (Dubey et al., 2024).
@@
 We directly download the best Zephyr-7B RMU model from the WMDP authors (https://huggingface.co/cais/Zephyr_RMU) for testing.
+For Mistral, we run a hyperparameter sweep over , layer indices 3,4,5, 4,5,6, and 5,6,7, and learning rates .
+We select runs with the lowest possible WMDP accuracies that don’t completely destroy MMLU accuracy.
+For Mistral, this is and lr= at layers 5,6,7.
+We sweep across the same hyperparameters for Llama-3-8B.
+Llama-3-8B-Instruct uses the best hyperparameters found in the base model sweep.
+The runs shown in Table 1 have and lr= at layers 4,5,6.
+All runs had a steering coefficient of 6.5.
 RepNoise (Rosati et al., 2024).
@@
 The method was run for one epoch with a batch size of 4.
+For Mistral, we run a hyperparameter sweep over , , and learning rates .
+We selected the run that has the lowest possible WMDP accuracies without destroying MMLU accuracy.
+This run, shown in Table 1, has the parameters , , lr=.
+We run a sweep over the same hyperparameters for Llama-3-8B, and use the best runs from the base model to decide hyperparameters for Llama-3-8B-Instruct. The runs shown in Table 1 had , , lr=.
 B.2 Harry Potter Results
@@
 This configuration strikes a balance between erasure efficacy, computational efficiency, and preservation of general language capabilities.
-Appendix E Conditional Fluency Training
+Appendix E Low-Rank vs Full Finetuning
+We analyze the role of using low-rank updates with ELM comparing its performance against finetuning the layers directly without any rank constraints.
+In Table 4, we show the performance of ELM on Zephyr-7B when editing with full finetuning and low-rank model editing.
+Full finetuning effects the specificity of the model and makes the unlearning broader damaging the general capabilities of the model.
+Where as low-rank model editing preserves the specificity while being effective at erasure.
+| Method | Innocence () | Specificity () | ||
+| Bio | Cyber | MMLU | MT-Bench | |
+| Original | 64.4 | 44.3 | 58.5 | 7.3 |
+| ELM - Full | 25.4 | 27.1 | 45.2 | 3.4 |
+| ELM - LoRA | 29.7 | 27.2 | 56.6 | 7.1 |
+Appendix F Conditional Fluency Training
 To achieve seamless editing, ELM must generate fluent text even when prompted about erased concepts.
@@
 This targeted approach ensures that we preserve the model’s general knowledge while specifically adapting its behavior for erased concepts.
-Appendix F Progression of ELM Training
+Appendix G Progression of ELM Training
 We evaluate the ELM intermediate checkpoints to observe the training dynamics of the method in Figure 5.
@@
 Bio-threat knowledge takes more time to be erased from the model - which could be directly proportional to the initial amount of prior knowledge.
+Appendix H Adversarial Attacks
+We also attack ELM with BEAST (Sadasivan et al., 2024), a fast adversarial prompt based attack on LLMs.
+BEAST finds an adversarial prompt that can be appended to the original attack prompt to generate target response.
+We find that BEAST is unable to extract erased information from ELM:
+Additionally - finetuning attack where we train ELM model autoregressively on the original forget dataset.
+We find that the resulting attacked model brings back the knowledge slightly (Bio: 29.7% to 42.2%; Cyber: 27.2% to 29.4%) but not to the original level of 64.4% Bio and 44.3% Cyber.
+ELM models can be retrained to bring back erased knowledge, but it is harder.

### v3

@@
 Abstract
-In this work, we propose Erasure of Language Memory (ELM), an approach to concept-level unlearning built on the principle of matching the distribution defined by an introspective classifier.
-Our key insight is that effective unlearning should leverage the model’s ability to evaluate its own knowledge, using the model itself as a classifier to identify and reduce the likelihood of generating content related to undesired concepts.
+In this work, we introduce Erasure of Language Memory (ELM), a principled approach to concept-level unlearning that operates by matching distributions defined by the model’s own introspective classification capabilities.
+Our key insight is that effective unlearning should leverage the model’s ability to evaluate its own knowledge, using the language model itself as a classifier to identify and reduce the likelihood of generating content related to undesired concepts.
 ELM applies this framework to create targeted low-rank updates that reduce generation probabilities for concept-specific content while preserving the model’s broader capabilities.
 We demonstrate ELM’s efficacy on biosecurity, cybersecurity, and literary domain erasure tasks.
-Comparative analysis shows that ELM achieves superior performance across key metrics, including near-random scores on erased topic assessments, maintained coherence in text generation, preserved accuracy on unrelated benchmarks, and robustness under adversarial attacks.
+Comparative evaluation reveals that ELM-modified models achieve near-random performance on assessments targeting erased concepts, while simultaneously preserving generation coherence, maintaining benchmark performance on unrelated tasks, and exhibiting strong robustness to adversarial attacks.
 Our code, data, and trained models are available at elm.baulab.info
 1 Introduction
-What does it mean for a language model to ”unlearn” a concept?
-While machine unlearning has traditionally focused on removing specific training samples from model memory, there is an increasing need to be able to erase broad conceptual knowledge—for example, removing all information about biological weapons rather than just a few training examples containing that information.
-In this paper we examine how concept-level unlearning leads to a new approach to knowledge removal in language models.
-Prior approaches to unlearning broadly fall into three categories: (1) retraining on filtered data (2) reversed-gradient-based methods that attempt to ”un-train” specific knowledge, and (3) representation manipulation approaches that disrupt internal activations for targeted content.
+What does it mean for a language model to "unlearn" a concept?
+While machine unlearning has traditionally focused on removing specific training samples from model memory, there is an increasing need to be able to erase broad conceptual knowledge—for example, removing all information about a dangerous concept like biological weapons.
+In this paper, we introduce a new approach to concept erasure in LLMs, which allows for seamless targeted removal of knowledge related to a particular broad concept by exploiting the model’s own ability to identify the undesired concept.
+Prior approaches to unlearning broadly fall into three categories: (1) retraining on filtered data (2) reversed-gradient-based methods that attempt to "un-train" specific knowledge, and (3) representation manipulation approaches that disrupt internal activations for targeted content.
 Unfortunately, each of these strategies have limitations that make them impractical for unlearning in large language models: dataset filtering requires retraining that is costly at scale; gradient reversal methods are unstable and create broad damage to the model; and representation manipulation creates obvious behavioral artifacts.
@@
 We propose a fundamentally different approach that leverages the model’s own ability to recognize and classify knowledge.
-Our key insight is that language models can act as their own critics: they can evaluate whether a piece of text demonstrates knowledge of a particular concept.
+Our key insight is that language models can act as their own critics: for any arbitrary piece of text, models can implicitly evaluate the probability of that text belonging to a particular concept.
 This self-classification provides a natural objective for unlearning: we can modify the model to reduce the likelihood of generating text it would classify as containing target concept.
 This insight leads to Erasure of Language Memory (ELM), a method that directly optimizes the model’s generation probabilities based on introspective classification.
-Unlike approaches like Representation Misdirection for Unlearning (RMU; Li et al., 2024) which manipulates internal activations without a clear behavioral target, or WhoIsHarryPotter (Eldan & Russinovich, 2023) which modifies training data but fails to fully eliminate concept knowledge, ELM has a principled objective: the model should generate coherent text that the language model itself would not classify as demonstrating knowledge of the target concept.
+Unlike approaches like Representation Misdirection for Unlearning (RMU; Li et al., 2024) which manipulates internal activations without a clear behavioral target, or WhoIsHarryPotter (Eldan and Russinovich, 2023) which develops heuristics for modifying training data that fail to fully eliminate concept knowledge (Section 5.7), ELM has a principled objective: the model should generate coherent text that the language model itself would not classify as demonstrating knowledge of the target concept.
 Our method achieves this through targeted fine-tuning that reduces the likelihood of generating content the model identifies as related to the concept being erased.
@@
 The self-classification framework also provides a natural way to ensure the model maintains coherent text generation even when prompted about erased concepts—it learns to generate alternative content that it would not classify as demonstrating the target knowledge.
-Prior unlearning methods face fundamental limitations when evaluated holistically.
-Our proposed evaluation desiderata—innocence (lack of target knowledge), specificity (preserved capabilities), and seamlessness (coherent generation)—reveal these tradeoffs: gradient reversal achieves innocence but degrades general capabilities, representation manipulation preserves capabilities but generates incoherent text, and dataset filtering maintains coherence but fails to fully eliminate knowledge.
-Through extensive experiments on WMDP biosecurity and cybersecurity benchmarks, as well as literary domain erasure, we measure the effectiveness of ELM’s principled approach to concept removal compared to previous methods.
-Our evaluation using both standard metrics and adversarial probing shows that ELM achieves robust concept erasure while maintaining model coherence and general capabilities.
+We compare our approach to prior methods, evaluating erased models under four desiderata: innocence (lack of target knowledge), specificity (preserved capabilities), seamlessness (coherent generation), and robustness to adversarial attacks.
+These criteria reveal tradeoffs: gradient reversal achieves innocence but degrades general capabilities, representation manipulation preserves capabilities but generates incoherent text, and dataset filtering maintains coherence but fails to fully eliminate knowledge.
+Through extensive experiments on WMDP biosecurity and cybersecurity benchmarks (Li et al., 2024), as well as literary domain erasure (Eldan and Russinovich, 2023), we we show that ELM achieves robust concept erasure while maintaining model coherence and general capabilities.
 2 Related work
 Machine Unlearning
-The idea of removing specific data from machine learning models, known as machine unlearning, has gained attention in recent years, initially motivated by privacy concerns (Cao & Yang, 2015; Harding et al., 2019).
+The idea of removing specific data from machine learning models, known as machine unlearning, has gained attention in recent years, initially motivated by privacy concerns (Cao and Yang, 2015; Harding et al., 2019).
 Early methods focused on efficiently removing individual training examples or facts from models (Golatkar et al., 2020; Ma et al., 2022; Jang et al., 2022a).
-However, most existing benchmarks evaluate unlearning on artificially created deletion sets (Choi & Na, 2023; Goel et al., 2022; Maini et al., 2024), in contrast to our focus on real-world distributions of broad conceptual knowledge.
+However, most existing benchmarks evaluate unlearning on artificially created deletion sets (Choi and Na, 2023; Goel et al., 2022; Maini et al., 2024), in contrast to our focus on real-world distributions of broad conceptual knowledge.
 Erasing broad conceptual knowledge from LLMs
-New approaches to machine unlearning have recently gained traction on the problem of removing dangerous capabilities from LLMs (Lynch et al., 2024; Ilharco et al., 2023; Jang et al., 2022b; Lu et al., 2022; Yu et al., 2023; Casper et al., 2024; Eldan & Russinovich, 2023).
-Our work is most directly comparable to three state-of-the-art techniques: Representation Misdirection for Unlearning (RMU) (Li et al., 2024), WhoIsHarryPotter (WHP) (Eldan & Russinovich, 2023), and Representation Noising (RepNoise) (Rosati et al., 2024).
-RMU fine-tunes models to align internal activations with random scaled vectors when processing targeted concepts.
-WHP (Eldan & Russinovich, 2023) employs a two-stage approach, first training a reinforced model for the concept being erased and then training an unlearned model to behave differently on the reinforced logits.
-RepNoise (Rosati et al., 2024) focuses on removing information about harmful representations across all layers of the LLM, by doing gradient ascent along with some representation noising (training internal representations to match Gaussian noise).
-While all these methods successfully reduce model performance on erased knowledge, our measurements reveal that these previous approaches fall short in one or more of the three erasing goals.
-In contrast, our current work erases concepts by fine-tuning towards a principled target distribution that is designed to balance all three desiderata of innocence, specificity, and seamlessness.
-Other methods such as LLMU (Yao et al., 2023), SSD (Foster et al., 2024), and SCRUB (Kurmanji et al., 2024) have also been proposed but face significant limitations.
-LLMU uses a differentiable membership classifier to identify unlearning set examples but struggles with imprecisely defined target distributions (see Li et al., 2024).
-SSD offers a retraining-free method for removing specific samples but is limited to individual examples rather than broader knowledge domains.
-SCRUB aims to remove sensitive data while preserving model performance on other tasks but requires access to the full training dataset, which is often impractical.
-Importantly, these methods were compared by RMU (Li et al., 2024) and found to be less effective for erasing broad conceptual knowledge.
+Recent machine unlearning approaches have addressed removing dangerous capabilities from LLMs (Lynch et al., 2024; Ilharco et al., 2023; Jang et al., 2022b; Lu et al., 2022; Yu et al., 2023; Casper et al., 2024; Eldan and Russinovich, 2023; Mekala et al., 2024).
+Our work directly compares with three state-of-the-art techniques: Representation Misdirection for Unlearning (RMU) (Li et al., 2024), which fine-tunes models to align internal activations with random scaled vectors when processing targeted concepts; WhoIsHarryPotter (WHP) (Eldan and Russinovich, 2023), which employs a two-stage approach with reinforced and unlearned models; and Representation Noising (RepNoise) (Rosati et al., 2024), which removes harmful representations via gradient ascent with representation noising.
+While these methods reduce model performance on erased knowledge, our measurements show they fall short in meeting all three erasing goals.
+Our work instead erases concepts by fine-tuning towards a principled target distribution designed to balance innocence, specificity, and seamlessness.
+Alternative methods including LLMU (Yao et al., 2023), SSD (Foster et al., 2024), and SCRUB (Kurmanji et al., 2024) face significant limitations: LLMU struggles with imprecisely defined target distributions (see Li et al., 2024); SSD only removes specific samples rather than broader knowledge domains; and SCRUB requires access to the full training dataset.
+Comparative analyses by RMU (Li et al., 2024) found these approaches less effective for erasing broad conceptual knowledge.
 Distilling generative model outputs.
-Some techniques for controlling generative model outputs involve using an auxiliary generative model to specify the desired behavior over some training data, and then distilling these outputs into a target model by training the target model to mimic the demonstrated behavior.
-For example Askell et al.
+Controlling generative model outputs often involves distillation: using auxiliary generative models to specify desired behavior, then training target models to mimic this behavior.
+Askell et al.
 (2021) and Bai et al.
-(2022) prompt an unsafe model into behaving more safely, and then distil the prompted model.
-Similarly, Gandikota et al.
-(2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes.
-(Rosati et al., 2024) can also be classified under distillation methods as it mimics a Gaussian distribution when processing harmful tokens.
-Similarly, ELM matches the harmful logits to a modified output distribution, but unlike the distillation methods that usually adhere to the standard matching objective, we employ a multi-objective framework to address the seamlessness and specificity concerns with standard distillation methods.
-Prior works like Emulator and DeRa leverage probability ratios for behavioral modification - Emulator (Mitchell et al., 2023) scales fine-tuning changes between model sizes, while DeRa (Liu et al., 2024) controls alignment strength during inference.
-In contrast, ELM introduces a simpler, principled approach focused specifically on reducing the likelihood of knowledge concept generation.
+(2022) prompt unsafe models into safer behavior before distillation, while Gandikota et al.
+(2023) train diffusion models to mimic edited versions that avoid generating certain attributes.
+Rosati et al.
+(2024) similarly mimics Gaussian distributions when processing harmful tokens.
+ELM also matches harmful logits to modified output distributions but employs a multi-objective framework addressing seamlessness and specificity concerns inherent in standard distillation.
+While prior works like Emulator (Mitchell et al., 2023) and DeRa (Liu et al., 2024) leverage probability ratios for behavioral modification, ELM introduces a simpler, principled approach specifically focused on reducing knowledge concept generation likelihood.
 Erasing in generative image models
@@
 (2024) erase concepts by modifying the key value mapping of cross attention layers in a low rank closed form update.
-Other works remove the knowledge of unwanted concepts from the model weights; proposing attention re-steering through fine-tuning (Zhang et al., 2023), fine-tuning the attention weights (Kumari et al., 2023) and continual learning (Heng & Soh, 2023).
-We take inspiration from (Gandikota et al., 2023) to reduce the likelihood of a concept being generated.
-3 Background
-3.1 Transformer
-We focus on autoregressive transformer-based models (Tunstall et al., 2023a; Dubey et al., 2024).
-Given an input token sequence , these models output a probability distribution over a vocabulary to predict the next token:
-The model architecture comprises stacked blocks, each usually containing a multi-head attention (MHA) layer followed by a multi-layer perceptron (MLP).
-These layers are stacked on top of each other, and their outputs are added to the residual stream of the model.
-MHA layers play an important role in bringing information from previous input tokens, while MLP layers are found to play the role of knowledge banks (Meng et al., 2022; Geva et al., 2023).
-These calculations can be summarized as follows,
-where denotes the hidden state for token at layer .
-Previous research (Meng et al., 2022) has localized much model knowledge within early to mid-layer blocks.
-As our own extensive experiments have similarly found good efficacy when erasing from earlier blocks of parameters, we restrict our parameter changes to MLP and MHA layers from those same blocks.
-3.2 Transformer Next Token Prediction: A Classification Perspective
-Language models are typically viewed through the lens of autoregressive sequence modeling, where the model predicts the next token given a sequence of previous tokens.
-However, these models can also be understood as powerful classifiers.
-To illustrate this connection, let’s examine both perspectives:
-Traditional Perspective: The standard way to describe an autoregressive language model is through the factorization:
+Other works remove the knowledge of unwanted concepts from the model weights; proposing attention re-steering through fine-tuning (Zhang et al., 2023), fine-tuning the attention weights (Kumari et al., 2023) and continual learning (Heng and Soh, 2023).
+We take inspiration from Gandikota et al.
+(2023) to reduce the likelihood of a concept being generated.
+3 Next Token Prediction: A Classification Perspective
+Language models are typically viewed through autoregressive sequence modeling, but they can also be understood as powerful text classifiers.
+The standard way to describe an autoregressive language model is:
 where the model predicts future tokens conditioned on previous tokens .
-Classification Perspective: We can reformulate this using Bayes’ rule to obtain:
-In this view, the model acts as a classifier where previous tokens serve as class labels for the text that follows.
-For instance, if we provide the context as ”The following text is about general biology”, the term represents the probability that the generated text belongs to the ”general biology” class.
-This classification perspective enables us to manipulate the model’s knowledge by adjusting these classification probabilities.
-We can define a target distribution that modifies how strongly the model associates generated text with specific concepts:
-where is a scaling factor that controls the strength of the modification. When , we increase the likelihood of generating text associated with the class ; when , we decrease it.
-To implement this in practice with an autoregressive language model, we apply Bayes’ rule again:
-In this form, both the numerator and denominator can be read out of the pretrained language model.
-Eq 6 is particularly powerful for concept unlearning because it allows us to use the language model itself as both the generator and classifier of content.
-For example, to unlearn bioweapons expertise, we can reduce the model’s probability of generating text that would naturally follow the context ”You are an expert in bioweapons.” This approach requires no external classifier - we leverage the model’s own understanding of concept relationships as encoded in its next-token predictions.
-Our perspective is inspired by the classifier-free guidance work (Ho & Salimans, 2022) in diffusion model research which has later been adopted to language models (Sanchez et al., 2023).
-3.3 Parameter-Efficient Fine-tuning via Low-Rank Adapters
-Fine-tuning large pre-trained language models involves large sets of parameters, which can result in overfitting.
-To address this issue, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021), a method that introduces a factorization of weights to significantly reduce the number of parameters.
-Consider a pre-trained model layer with weights , where and denote input and output dimensions, respectively.
-LoRA decomposes the weight update as follows:
-Here, represents a small rank (), constraining the update to a low-dimensional subspace.
-By optimizing only and while keeping fixed, LoRA significantly reduces the trainable parameters.
-During inference, the update can be efficiently incorporated into the weights:
-The LoRA approach offers a balance between adaptation capability, reduced overfitting, and parameter efficiency, making it particularly suitable for our setting of fine-tuning large-scale models.
+Classification Perspective.
+We can also think of previous tokens as a “class label” for whatever arbitrary document follows those tokens.
+For example, say that a prefix consists of the tokens “Here is a text about biology.” Conditioned on that prefix, we would expect a news article about finance to have a much lower probability than a chapter from a biology textbook.
+To reflect this intuition, we can rewrite the previous equation using Bayes’ Rule:
+and specifically interpret as the probability that a piece of text belongs to the “biology” class.
+This perspective enables us to manipulate the model’s output by adjusting these classification probabilities for a particular prefix using a scaling parameter :
+where controls the likelihood of a text belonging to .
+When , we increase the likelihood of generating text associated with the class ; when , we decrease it.
+For implementation in an autoregressive language model, we apply Bayes’ rule again:
 4 Method
@@
 While traditional unlearning methods have focused on sample removal through dataset retraining, gradient ascent, or representation disruption, ELM leverages the language model’s own ability to evaluate and modify its knowledge.
-Specifically, ELM uses the language model itself as its own introspective classifier, as established in Section 3.2.
-Rather than training external classifiers or explicitly manipulating representations, we leverage the model’s next-token predictions to identify and reduce the likelihood of generating concept-specific content.
-This allows us to operate directly on the model’s probability distributions, providing a principled approach to unlearning.
-The core of our method is a self-classification objective that reduces the likelihood of generating text that the model would classify as containing the target concept.
-Given a dataset containing text related to the target concept, we modify the model’s predicted probabilities to diverge from generating content related to the concept being erased and instead favor generating content related to a safer alternative :
-where controls the strength of knowledge modification. is the concept to be erased, and is an alternative concept. The corresponding loss is:
-In practice, we encounter two main challenges when implementing this objective.
-First, knowledge in language models is often entangled - modifying one concept can unintentionally affect related concepts.
-To address this, we preserve the model’s behavior on a set of related but safe concepts by matching its original distribution on retain data, , containing unrelated text:
-Second, the self-classification objective alone might lead to incoherent text generation when prompted about erased concepts.
-We address this by applying our core objective (Equation 4) during inference to generate synthetic training examples, then using these to train the model in an autoregressive setting to maintain coherent generation:
-The final training objective combines these terms with appropriate weights:
-We implement this through low-rank adapters attached to early model layers, allowing precise modification of the model’s knowledge while maintaining its broader capabilities.
-This approach provides a principled method for concept unlearning that avoids the instability of gradient reversal methods and the incoherence of representation hacking approaches.
+Specifically, we leverage the implicit classification behavior of the language model (Section 3) as a training signal, and use this signal to train targeted low-rank adapters on a subset of model layers.
+4.1 Concept Unlearning via Self Classification
+The core of our method is a self-classification objective that reduces the likelihood of generating text the model would classify as containing the target concept.
+We work with an erase dataset containing text sequences related to the concept we want to forget.
+To implement our approach, we use two context prompts: representing the concept to be erased (e.g., "This text is written by a specialist in bioweapons"), and representing an alternative distribution (e.g., "This text is written by a novice with no knowledge of bioweapons").
+Our goal is to inhibit the model’s internal classifier: when processing dangerous documents , we want the model’s internal classifier to look more like it does in the setting and less like it does for the concept .
+In other words, when the erased model encounters a dangerous input prompt, it should behave more like a “novice” and less like an “expert”:
+where represents the probability distribution from the original pre-trained model with parameters , controls the strength of knowledge modification, and the rightmost term comes from Equation 4.
+We can frame this in terms of next-token prediction as follows:
+where the intuition is that key tokens that are more likely to be output when prefixed by are promoted, whereas tokens that are more likely to be output under are quashed (see Figure 1 for an example).
+The corresponding loss function compares this classifier modified distribution to the distribution of the ELM model with parameters :
+In practice, we encounter a significant challenge when implementing our objective: knowledge in language models is often entangled - modifying one concept can unintentionally affect related concepts.
+To address this, we preserve the model’s behavior on a set of related but safe concepts by using a retention dataset containing text sequences unrelated to the erased concept.
+We train the model to match its original distribution on this data:
+4.2 Optional Fluency Enhancement for Smaller Models
+For smaller models, we observe that the self-classification objective alone might lead to incoherent text generation when prompted about erased concepts.
+To maintain natural text generation in these cases, we apply our core objective (Equation 5) during inference to generate synthetic training examples.
+For each prompt from , we generate tokens using our probability modifier and train the model in an autoregressive setting on the generated tokens to maintain coherence.
+More details are provided in Appendix E:
+When this additional fluency term is included, the total loss becomes:
+This provides a principled method for concept unlearning that avoids the instability of gradient reversal methods and the incoherence of representation hacking approaches.
+4.3 Low-Rank Adapters
+Previous research (Meng et al., 2022; Geva et al., 2023) has localized model knowledge within early to mid-layer blocks.
+We find that low-rank adapters (Hu et al., 2021) trained on early layers allow for the most precise modification of model knowledge while maintaining broader capabilities.
+Compared to general fine-tuning, low-rank adapters allow for targeted unlearning, without damaging unrelated knowledge (Appendix D.2).
+Consistent with previous work, we find that these adapters are most effective at early layers (Figure 4).
 5 Experiments
@@
 Models.
-We apply ELM to a range of state-of-the-art language models, including Zephyr-7B Beta (Tunstall et al., 2023b), Mistral-7B (Jiang et al., 2023), Llama3-8B, and Llama3-8B-instruct (Dubey et al., 2024) for the WMDP erasure tasks.
-For the Harry Potter knowledge erasure, we use the Llama-2-7B Chat model (Touvron et al., 2023) to maintain consistency with prior work from (Eldan & Russinovich, 2023).
+We apply ELM to a range of state-of-the-art language models, including Zephyr-7B Beta (Tunstall et al., 2023), Mistral-7B (Jiang et al., 2023), Llama3-8B, LLama3-70B, Llama3-8B-instruct (Dubey et al., 2024), and Qwen2.5-32B (Yang et al., 2024) for the WMDP erasure tasks.
+For the Harry Potter knowledge erasure, we use the Llama-2-7B Chat model (Touvron et al., 2023) to maintain consistency with prior work from Eldan and Russinovich (2023).
 This selection of models enables us to evaluate ELM’s performance across various model architectures and training paradigms.
@@
 To facilitate conditional erasure (Eq.
-4), we prepend contexts such as “You are an expert in” followed by concept-specific keywords.
-Additionally, we incorporate text completion examples for consistency, following the approach used by (Qi et al., 2024).
+5), we prepend contexts such as “You are an expert in” followed by concept-specific keywords.
+Additionally, we incorporate text completion examples for consistency, following the approach used by Qi et al.
+(2024).
+We show more details in Appendix C
 Evaluation Metrics.
-We assess our method, Erasure of Language Memory (ELM), along four key dimensions:
+We assess our method, Erasure of Language Memory (ELM), along four key dimensions. We provide implementation details in Appendix B:
 1.
@@
 […]
-| RMU | 33.5 | 28.7 | 27.1 | 1.0 | 29.9 | |
-| RepNoise | 35.3 | 39.6 | 55.0 | 2.1 | 26.7 | |
-| Ours | 28.7 | 26.4 | 55.4 | 3.7 | 15.3 | |
+| RMU | 49.4 | 37.0 | 40.1 | 3.9 | 4.1 | |
+| RepNoise | 54.7 | 43.6 | 54.2 | 5.5 | 4.9 | |
+| Ours | 33.3 | 26.6 | 57.2 | 4.8 | 4.5 | |
 […]
-| RMU | 66.8 | 45.8 | 56.5 | 7.4 | 3.0 | |
+| RMU | 46.2 | 31.9 | 56.5 | 7.4 | 3.0 | |
 | RepNoise | 59.9 | 44.1 | 60.1 | 6.7 | 3.5 | |
@@
 […]
-| RMU | 49.4 | 37.0 | 40.1 | 3.9 | 4.1 | |
-| RepNoise | 54.7 | 43.6 | 54.2 | 5.5 | 4.9 | |
-| Ours | 33.3 | 26.6 | 57.2 | 4.8 | 4.5 | |
+| Ours | 33.1 | 27.1 | 78.4 | 7.9 | 4.8 | |
+| Ours () | 32.7 | 27.5 | 78.8 | 7.8 | 5.1 | |
+[…]
+| Ours | 33.7 | 28.2 | 75.2 | 7.2 | 4.8 | |
+| Ours () | 32.1 | 28.0 | 75.7 | 7.2 | 4.3 | |
 As shown in Table 1, ELM consistently achieves near-random performance (random guess is 25%) on erased WMDP concepts (Bio and Cyber) while maintaining high scores on general knowledge (MMLU) and language understanding (MT-Bench) tasks.
 Notably, ELM demonstrates superior fluency when generating text related to erased concepts, as evidenced by lower reverse perplexity scores compared to RMU and RepNoise.
+We observe an emergent artifact in larger models where the fluency term () is no longer necessary for maintaining fluency.
+Our erasing loss term is enough to precisely erase the knowledge while maintaining fluency at the same time.
+This suggests that larger models have more precise internal classifiers than smaller models, making additional fluency objectives unnecessary.
+We show the progression on unlearning in Appendix F and qualitative samples in Appendix H
 5.3 Ablation Study
@@
 Table 2 shows the impact on concept erasure (WMDP), general knowledge (MMLU), and generation quality (MT-Bench, Perplexity) when removing or modifying individual terms.
+We show more fine-grained ablations and hyper-parameters in Appendix D
 First, proves crucial for innocence.
@@
 7 with random vector) can achieve some erasure in cyber concepts, they lead to incoherent outputs.
-The retain term (Eq. 11) is vital for specificity. Its removal yields the lowest MMLU scores, demonstrating its role in maintaining broad knowledge while enabling targeted erasure.
+The retain term (Eq.
+8) is vital for specificity.
+Its removal yields the lowest MMLU scores, demonstrating its role in maintaining broad knowledge while enabling targeted erasure.
 Ablating , leads to effective erasure but the model generates low-quality gibberish text when prompted for erased concept.
-Replacing it with random text from WikiText (Merity et al., 2016) slightly reduces fluency while maintaining erasure effectiveness.
+Replacing it with random text from WikiText Merity et al.
+(2016) slightly reduces fluency while maintaining erasure effectiveness.
 Qualitatively this produces awkward outputs that tend to be irrelevant to input prompts.
 This underscores the term’s role in maintaining seamless contextual relevance.
+However, we note that is not necessary for larger models (Qwen2.5-32B and Llama3-70B).
+The full ELM method achieves the best balance between concept erasure and general performance. We show a qualitative example with each of the settings below:
 | Setup | Innocence () | Specificity () | Seamlessness () | ||
@@
 […]
-The full ELM method achieves the best balance between concept erasure and general performance. We show a qualitative example with each of the settings below:
 5.4 Specificity Analysis
@@
 5.5 Robustness to Attacks
-We evaluate the robustness of ELM by employing Greedy Coordinate Gradient attack (GCG) (Zou et al., 2023).
+We evaluate the robustness of ELM by employing Greedy Coordinate Gradient attack (GCG) Zou et al.
+(2023).
 We optimize for an adversarial suffix prompt to induce the model to generate a target text.
@@
 ELM demonstrates strong resistance to these attacks even when optimized for 5000 steps, whereas, the original model starts responding to the query with the GCG suffix, as illustrated in the example below.
-We provide GCG attack details in Appendix C.
-We also show ELM’s robustness to more attacks in Appendix H.
+We provide more details on GCG attack and show other adversarial attacks in Appendix G.
 5.6 Probing and Activation Analysis
-To estimate the presence of erased knowledge within the internal representations of a model, we conduct the probing analysis, training a linear probe using the same setup as used by (Li et al., 2024).
+To estimate the presence of erased knowledge within the internal representations of a model, we conduct the probing analysis, training a linear probe using the same setup as used by Li et al. (2024).
 The results in Figure 3(a) reveal distinct knowledge retention patterns across methods.
 ELM and RMU achieve effective erasure, maintaining low probe accuracies across all layers for both biosecurity and cybersecurity MCQs.
-In contrast, RepNoise shows partial retention, particularly for cybersecurity concepts.
+In contrast, RepNoise shows partial retention, particularly for WMDP-Cyber.
 Analysis of activation norms, in Figure 3(b), further highlights the differences.
@@
 To further demonstrate the versatility of ELM, we apply it to the unlearn knowledge of Harry Potter literary universe.
-We compare ELM against RMU and WhoIsHarryPotter (WHP) (Eldan & Russinovich, 2023) methods for Llama-2-7B Chat.
+We compare ELM against RMU and WhoIsHarryPotter (WHP) (Eldan and Russinovich, 2023) methods for Llama-2-7B Chat.
 Table 3 presents this comparison.
 ELM achieves a balance between effective knowledge erasure (low HP MCQ score) and maintaining fluent generation (low reverse-perplexity).
-Similar to (Lynch et al., 2024), we found WHP model (Eldan & Russinovich, 2023) maintains fluency but fails to effectively erase the target knowledge.
+Similar to Lynch et al.
+(2024), we found WHP model (Eldan and Russinovich, 2023) maintains fluency but fails to effectively erase the target knowledge as revealed in its retained ability to answer multiple-choice questions about Harry Potter.
 RMU (Li et al., 2024) proved to be ineffective in erasing with a large hyper parameter sweep.
@@
 The method shows some degradation in performance on semantically adjacent concepts, indicating that our approach may need refinement to achieve more precise boundaries between related knowledge.
-Additionally, although generated text maintains basic fluency, it sometimes lacks semantic coherence - suggesting that our probability modification may be overly aggressive in some cases.
+Additionally, although generated text maintains basic fluency, it sometimes lacks semantic coherence, suggesting that our probability modification may be overly aggressive in some cases.
 The most significant challenge lies in handling deeply interconnected concepts, where modifying the model’s behavior for one concept may have ripple effects through its broader knowledge base.
@@
 ELM is available as open-source code. Source code, trained models, and data sets for reproducing our results can be found at elm.baulab.info and at https://github.com/rohitgandikota/erasing-llm.
-Appendix A Details on metrics
+Appendix A Impact Statement
+In this work, we develop a framework for thinking about concept erasure in language models, as well as a new approach to erasing conceptual knowledge.
+Although we focus on removal of potentially harmful knowledge, this technology could be misused to remove legitimate knowledge from a language model without users’ awareness.
+Additionally, if our method is used to remove harmful knowledge, it may create a false sense of security, as models could retain harmful knowledge that is undetected by our metrics.
+Unlearning has an important place in safety considerations for language models, but should not be the only approach.
+Finally, we also acknowledge that our evaluations are focused on harmful knowledge encoded in English; we have not evaluated this approach cross-linguistically.
+We release our code publicly to enable open and safe research.
+Appendix B Details on metrics
 Multiple Choice Questions.
@@
 To measure the seamlessness of edits, we aim to quantify the fluency of the text being generated by the edited model when prompted with the concept being erased.
-To evaluate this we prompt the models using questions from MCQ dataset from WMDP (Li et al., 2024) and let the models generate text free-form up to 500 tokens.
+To evaluate this we prompt the models using questions from MCQ dataset from WMDP Li et al.
+(2024) and let the models generate text free-form up to 500 tokens.
 We then measure the perplexity on generated text using a totally different evaluation model, Llama3.1-8B (Dubey et al., 2024).
-Appendix B Baseline Methods
-B.1 WMDP Results
+Appendix C Baseline Methods
+We compare ELM against other baselines across different models for unlearning WMDP-Bio and WMPD-cyber in Table 4.
+ELM shows stronger general erasure performance across different model architectures and settings.
+C.1 WMDP Results
 RMU (Li et al., 2024).
@@
 We run a sweep over the same hyperparameters for Llama-3-8B, and use the best runs from the base model to decide hyperparameters for Llama-3-8B-Instruct. The runs shown in Table 1 had , , lr=.
-B.2 Harry Potter Results
+| Model | Method | Innocence () | Specificity () | Seamlessness | ||
+| Bio | Cyber | MMLU | MT-Bench | R-PPL () | ||
+| Zephyr-7B | Original | 64.4 | 44.3 | 58.5 | 7.3 | 6.0 |
+| RMU | 30.5 | 27.3 | 57.5 | 7.2 | 24.8 | |
+| RepNoise | 29.7 | 37.7 | 53.3 | 6.6 | 25.0 | |
+| Ours | 29.7 | 27.2 | 56.6 | 7.1 | 10.9 | |
+[…]
+| RMU | 33.5 | 28.7 | 27.1 | 1.0 | 29.9 | |
+| RepNoise | 35.3 | 39.6 | 55.0 | 2.1 | 26.7 | |
+| Ours | 28.7 | 26.4 | 55.4 | 3.7 | 15.3 | |
+[…]
+| RMU | 46.2 | 31.9 | 56.5 | 7.4 | 3.0 | |
+| RepNoise | 59.9 | 44.1 | 60.1 | 6.7 | 3.5 | |
+| Ours | 32.2 | 27.2 | 61.6 | 7.7 | 7.4 | |
+[…]
+| RMU | 49.4 | 37.0 | 40.1 | 3.9 | 4.1 | |
+| RepNoise | 54.7 | 43.6 | 54.2 | 5.5 | 4.9 | |
+| Ours | 33.3 | 26.6 | 57.2 | 4.8 | 4.5 | |
+[…]
+| Ours | 33.1 | 27.1 | 78.4 | 7.9 | 4.8 | |
+| Ours () | 32.7 | 27.5 | 78.8 | 7.8 | 5.1 | |
+[…]
+| Ours | 33.7 | 28.2 | 75.2 | 7.2 | 4.8 | |
+| Ours () | 32.1 | 28.0 | 75.7 | 7.2 | 4.3 | |
+C.2 Harry Potter Results
 RMU (Li et al., 2024).
@@
 The Harry Potter dataset used for RMU was not the exact same dataset used for ELM (https://huggingface.co/datasets/mickume/harry_potter_tiny), as performance was much worse for RMU on the latter dataset.
-WHP (Eldan & Russinovich, 2023).
+WHP (Eldan and Russinovich, 2023).
 We directly download the best Llama-2-7B Chat model from the original authors (https://huggingface.co/microsoft/Llama2-7b-WhoIsHarryPotter).
-Appendix C Robustness Evaluation
-To evaluate the robustness of ELM against adversarial attacks, we employ the Greedy Coordinate Gradient (GCG) method (Zou et al., 2023), utilizing the standard implementation from GraySwanAI (GraySwanAI, 2024).
-The GCG attack requires defining an initial prompt, a multi-token target text, and an initialized adversarial suffix.
-Following the protocol established in Li et al.
-(2024), we use a 20-token adversarial suffix and derive prompts from the WMDP MCQ datasets.
-To facilitate open-ended generation, we present only the question component of these prompts, omitting the multiple-choice structure.
-Our experiments reveal a stark contrast in robustness between ELM models and their base model counterparts.
-Even after extensive optimization exceeding 5000 iterations, we fail to identify a GCG prompt capable of inducing ELM models to generate content related to erased concepts.
-This resilience stands in marked contrast to the original models, which succumb to effective attack suffixes within 200 iterations, subsequently producing potentially harmful text.
-ELM:
-RMU:
-RepNoise:
 Appendix D Hyperparameter Analysis
@@
 This configuration strikes a balance between erasure efficacy, computational efficiency, and preservation of general language capabilities.
-Appendix E Low-Rank vs Full Finetuning
+D.1 Ablation on ELM Loss Terms
+We sweep the values of , , from ELM Loss terms in Equation 10.
+We run this ablation on Llama3-8B model and show the results in Table 5.
+We find that increasing the erase loss scale () tends to increase the erasure effect.
+Increasing the retain loss term () improves the specificity of the erasure.
+Finally, increasing the consistency term has improved fluency, but increasing it beyond a certain value affects the erasure efficacy of the method.
+| Value | WMDP-Bio () | MMLU () | R-PPL () | |
+| 0.0 | 0.70 | 0.63 | 7.13 | |
+| 0.5 | 0.43 | 0.62 | 12.49 | |
+| 1.0 | 0.37 | 0.62 | 7.92 | |
+| 1.5 | 0.34 | 0.62 | 9.79 | |
+| 2.0 | 0.35 | 0.62 | 8.80 | |
+| 0.0 | 0.25 | 0.24 | 9.19 | |
+| 0.5 | 0.31 | 0.61 | 10.41 | |
+| 1.0 | 0.37 | 0.62 | 7.92 | |
+| 1.5 | 0.38 | 0.62 | 10.72 | |
+| 2.0 | 0.37 | 0.62 | 9.31 | |
+| 0.0 | 0.28 | 0.61 | 22.29 | |
+| 0.5 | 0.34 | 0.62 | 8.34 | |
+| 1.0 | 0.37 | 0.62 | 7.92 | |
+| 1.5 | 0.39 | 0.62 | 12.01 | |
+| 2.0 | 0.35 | 0.62 | 11.64 |
+D.2 Low-Rank vs Full Finetuning
 We analyze the role of using low-rank updates with ELM comparing its performance against finetuning the layers directly without any rank constraints.
@@
 Full finetuning effects the specificity of the model and makes the unlearning broader damaging the general capabilities of the model.
-Where as low-rank model editing preserves the specificity while being effective at erasure.
+Low-rank model editing preserves the specificity while being effective at erasure.
 | Method | Innocence () | Specificity () | ||
@@
 | ELM - LoRA | 29.7 | 27.2 | 56.6 | 7.1 |
-Appendix F Conditional Fluency Training
-To achieve seamless editing, ELM must generate fluent text even when prompted about erased concepts.
+Appendix E Conditional Fluency Training
+For smaller models, we find that erasure loss alone is not enough to maintain fluency.
+To achieve seamless editing for smaller models, ELM must generate fluent text even when prompted about erased concepts.
 The ideal behavior mimics a model that never encountered the concept during pretraining.
@@
 This targeted approach ensures that we preserve the model’s general knowledge while specifically adapting its behavior for erased concepts.
-Appendix G Progression of ELM Training
+Appendix F Progression of ELM Training
 We evaluate the ELM intermediate checkpoints to observe the training dynamics of the method in Figure 5.
@@
 Bio-threat knowledge takes more time to be erased from the model - which could be directly proportional to the initial amount of prior knowledge.
-Appendix H Adversarial Attacks
-We also attack ELM with BEAST (Sadasivan et al., 2024), a fast adversarial prompt based attack on LLMs.
+Appendix G Robustness Evaluation
+G.1 Greedy Coordinate Gradient (GCG)
+To evaluate the robustness of ELM against adversarial attacks, we employ the Greedy Coordinate Gradient (GCG) method (Zou et al., 2023), utilizing the standard implementation from GraySwanAI (GraySwanAI, 2024).
+The GCG attack requires defining an initial prompt, a multi-token target text, and an initialized adversarial suffix.
+Following the protocol established in Li et al.
+(2024), we use a 20-token adversarial suffix and derive prompts from the WMDP MCQ datasets.
+To facilitate open-ended generation, we present only the question component of these prompts, omitting the multiple-choice structure.
+Our experiments reveal a stark contrast in robustness between ELM models and their base model counterparts.
+Even after extensive optimization exceeding 5000 iterations, we fail to identify a GCG prompt capable of inducing ELM models to generate content related to erased concepts.
+This resilience stands in marked contrast to the original models, which succumb to effective attack suffixes within 200 iterations, subsequently producing potentially harmful text.
+ELM:
+RMU:
+RepNoise:
+G.2 BEAST
+We also attack ELM with BEAST Sadasivan et al.
+(2024), a fast adversarial prompt based attack on LLMs.
 BEAST finds an adversarial prompt that can be appended to the original attack prompt to generate target response.
 We find that BEAST is unable to extract erased information from ELM:
+G.3 Finetuning Attack
 Additionally - finetuning attack where we train ELM model autoregressively on the original forget dataset.